### v1

# Parallel transport on matrix manifolds and Exponential Action

Abstract.
We express parallel transport for several common matrix Lie groups with a family of pseudo-Riemannian metrics in terms of matrix exponential and exponential actions. The expression for parallel transport is preserved by taking the quotient under certain scenarios. In particular, for a Stiefel manifold of orthogonal matrices of size $n\times d$, we give an expression for parallel transport along a geodesic from time zero to $t$, that could be computed with time complexity of $O(nd^{2})$ for small $t$, and of $O(td^{3})$ for large $t$, contributing a step in a long-standing open problem in matrix manifolds. A similar result holds for flag manifolds with the canonical metric. We also show the parallel transport formulas for the generalized linear group, and the special orthogonal group under these metrics.
Key words and phrases:
Parallel transport, Stiefel manifolds, Flag manifolds, Matrix Exponential, Matrix manifolds, Reductive Group.
1991 Mathematics Subject Classification:
15A16, 15A18, 15B10, 22E70, 51F25, 53C80, 53Z99
1. Introduction
In recent years, Riemannian geometry has found many applications in optimization, image, vision, and in modeling data with constraints and symmetry. A fundamental concept in geometry is parallel transport. While a tangent vector in Euclidean space can be transported between different points with almost no additional computation, moving a tangent vector between points in Riemannian geometry is a challenging yet fundamental problem. Parallel transport allows one to move tangent vectors along curves in a manifold in a way consistent with geometry, which is important in clustering problems, and in algorithms that require accounting for past increments. For example, the well-known optimization methods L-BFGS or conjugate gradient on an Euclidean space keep a history of past movements and use them to decide the next moves. These increments are represented as tangent vectors at different points in the iteration and must be transported to the current iterative point for aggregation. Similar issues arise in computer vision.
In the now classical paper [5] twenty-five years ago, the authors twice mentioned that there was no known $O(nd^{2})$ algorithm for parallel transport on the Stiefel manifold $\mathrm{St}_{n,d}$. To our best knowledge, this problem remains unsolved. In this paper, we contribute toward a solution for this problem with the help of exponential action. We show with an appropriate transformation, a class of metrics on matrix manifolds have geodesics computable by matrix exponential, while the parallel transport equation could be reduced to a linear ODE with constant coefficients, and thus could be solved by exponential action. For the case of the Stiefel manifolds, we show any tangent vector could be split into two parts, one in a submanifold similar to the great circle, where we can use exponential action. The other part, normal to a vector subspace containing the geodesic, can be transported using matrix exponential. We establish a similar result for flag manifolds with the canonical metric. We also show that the formula for the Stiefel manifold is a special case for pseudo-Riemannian metrics in a family with closed-form geodesic and parallel transport. Besides the state-of-the-art algorithm in the Stiefel case, we hope our results will provide insight into the properties of parallel transport.
1.1. Contribution
We give an explicit, with cost $O(C_{1}nd^{2}+C_{2}td^{3})$ for the parallel transport of Stiefel manifolds. Our formula for the parallel transport of a tangent vector $\eta$ stating at a point $Y\in\mathrm{St}_{n,d}$, the Stiefel manifold of orthogonal matrices in $\mathbb{R}^{n\times d}$ equipped with the metric eq. 44 defined by a parameter $\alpha>0$ (the canonical metric corresponds to $\alpha=\frac{1}{2}$ and the Euclidean embedded metric corresponds to $\alpha=1$) along a geodesic with initial speed $\xi$ is given by
(1) $\begin{gathered}\Delta(t)=\begin{bmatrix}Y|Q\end{bmatrix}\operatorname{expm}(t\begin{bmatrix}2\alpha A&-R^{\mathsf{T}}\\ R&0\end{bmatrix})\operatorname{expv}(t\mathrm{P},\begin{bmatrix}Y|Q\end{bmatrix}^{\mathsf{T}}\eta)\operatorname{expm}(t(1-2\alpha)A)\\ +(\eta-YY^{\mathsf{T}}\eta-QQ^{\mathsf{T}}\eta)\operatorname{expm}(t(1-\alpha)A)\end{gathered}$
where $Q$ is an orthogonal basis of the column vectors of $\eta-YY^{\mathsf{T}}\eta$, $A$ and $R$ are such that $\xi=YA+QR$ and $\mathrm{P}$ is an operator defined in eq. 52, the column-wise concatenation of $Y$ and $Q$ is denoted $\begin{bmatrix}Y|Q\end{bmatrix}$, and $\operatorname{expv}$ is the exponential action. The formula can be evaluated with cost $O(C_{1}nd^{2}+C_{2}td^{3})$ with $n>d$ because $Q,A,R$ could be computed with cost $O(nd^{2})$ [5, Corollary 2.2], the matrix exponentials is of cost $O(d^{3})$ (or more precisely $O(d^{3}\log_{2}t))$, the exponential action is of cost $O(td^{3})$ and the matrix multiplications are $O(nd^{2})$. Thus, the time complexity is $O(C_{1}nd^{2}+C_{2}td^{3})$. For small $t$, which is the situation in optimization problems, the first $O(nd^{2})$ term dominates the cost, whereas, for large $t$, the second $O(td^{3})$ will dominate.
The framework to derive this formula is also applicable to manifolds related to Stiefel manifolds, giving us a new parallel transport formula for flag manifolds with the canonical metric and recovering the known formula for Grassmann manifolds.
We also show the expression of parallel transport in terms of exponential action is applicable to a larger class of matrix groups and certain quotient spaces. We denote the relevant subspaces transposable subspaces, a property that in the Lie group literature is denoted reductivity but that we here use in the matrix context to obtain closed-form geodesics and parallel transport. With this result, we show explicit parallel transport formulas for this family of metrics for the generalized linear groups, and the special orthogonal groups. While not discussed here, the symplectic Lie algebra [6] is transposable, thus the results here also apply to symplectic groups.
1.2. Related literature and State-of-the-Art
Closed-form geodesics formulas for the generalized linear group are in [28, 18, 3]. In the original paper [5], Edelman et. al. gave an equation (equation 2.27) for parallel transport for the Stiefel manifold with the canonical metric. The differential equation is linear but even with exponential action, the cost is $O(n^{3})$. They also mentioned that this formula is a special case for reductive spaces (in particular, compact groups) and also gave an $O(nd^{2})$ algorithm for the Grassmann transport. As mentioned, reductivity overlaps, if not identical with our notion of transposability, but the latter notion is easy to formulate and leads quickly to explicit and computationally efficient formulas. We avoid the $O(n^{3})$ cost by exploiting a special property of Stiefel and flag manifolds, including Grassmann manifolds.
In [12], the authors give a formula for parallel transport for Stiefel manifolds in the Euclidean embedded metric ($\alpha=1$ in eq. 1). The formula is in terms of a function $F$ satisfying a matrix ODE, depending on the equation of the curve, and requires vectorization. The geodesic is calculated with $O(n^{3})$ cost for a given time. Other than that, we are not aware of progress in parallel transport for Stiefel manifolds. Numerically, in [9, 16, 15], the authors introduce numerical algorithms to solve transport problems, in particular the ladder schemes. As closed-form formulas are not available most of the time, they are of great importance.
The family of metrics considered here could be considered a deformation of a pseudo-Riemannian metric (also called semi-Riemannian metric) to a Riemannian metric, see [8], the deformation is often called the Cheeger deformation, although it has also been used by other authors. The Cheeger deformation comes from a (pseudo-Riemannian) submersion, thus, in geometry, is used to produce metrics with desired curvature properties. Geodesics could be lifted in a submersion.
In [13], the authors analyze sub-Riemannian geodesic arising from this family and note that this type of deformation produces closed-form geodesic. The resulting induced metrics for Stiefel manifolds are studied in [11]. In [20], one of the authors compute the connection and curvature for this family in the case of a bi-invariant metric, and proved proposition 1 in that case. Flag manifolds are studied by many authors, where Riemannian Hessian, exponential map, logarithm and the Levi-Civita connection are studied in [30, 24, 27, 21, 19].
The fact that parallel transport is an isometry is of interest in the optimization literature, allowing estimates for convergence analysis, see [26, 25].
1.3. Outline
In the next section, we review the background required in the paper, including bilinear form and pseudo-Riemannian geometry. We introduce the concept of transposable Lie-subalgebra, which includes $\mathfrak{gl}$ and $\mathfrak{so}$. We introduce the family of tractable metrics in theorem 1 and discuss the condition for the metrics to be Riemmanian.
We show the geodesic and parallel transport formula for generalized linear groups in section 4 and special orthogonal groups in section 5. In section 6, we show the quotient by a subgroup of particular structures will give closed-form geodesics. The parallel transport equation can be reduced to a linear ODE with variable coefficients in the general case, and constant coefficients, solvable by exponential action in special cases. In section 7, we provide parallel transport formulas for the Stiefel and flag manifolds, with a discussion of Grassmann manifolds in appendix A. In section 9, we provide numerical experiments. We end the paper with a discussion.
2. Background Review
2.1. Exponential Action
The solution of a linear ODE with constant coefficient $\dot{X}=AX$ and initial condition $X(0)=X_{0}$ is given by $\operatorname{expm}(tA)X_{0}$, where typically $X$ is a vector and $A$ is a matrix, where $\operatorname{expm}$ is the matrix exponential. We often need to treat the case where $X_{0}$ is a matrix and $A$ is an operator on a matrix space. In that situation, which is the situation encountered in this article, computing the exponential by vectorization is usually not efficient. The better alternative is to compute the exponential action, denoted by $\operatorname{expv}(tA,X_{0})$. Instead of evaluating the full matrix exponential, we approximate the action $\operatorname{expm}(tA)X_{0}$ more efficiently, often [2] by evaluating a scaling and squaring of a Taylor expansion of the operator exponential operating on $X_{0}$.
Specifically, the popular library algorithm for exponential action in Python (function $expm{\_}multiply$ in the scipy package [29] and similar functions for other languages, $expmv$ in MATLAB and $expv$ in Julia) are based on the work of Al-Mohy and Higham [2], approximating $e^{A}B\approx(T_{m}(s^{-1}A))^{s}B$, where $T_{m}$ denotes the $m$ terms Taylor series expansion of $e^{A}$. The cited paper provides an algorithm to find a pair of optimal $(m,s)$, based on estimates of $\|A^{p}\|^{1/p}_{1}$ for several values of $p$. The paper found the higher $p$ estimates are beneficial when $A$ is not normal, for example, when $A$ is upper triangular. These estimates are used to determine the Taylor order $m$.
The cost is thus a multiple of the cost of evaluating $Av$ for a vector (or matrix) $v$ of the same size with $X_{0}$. The multiplier is dependent on the required accuracy and $|t|\|A\|$, for an appropriate norm $\|A\|$. In library implementations, $A$ could be given as a function. To estimate the norm, we also need to evaluate the transpose $A^{\mathsf{T}}$.
In these implementations, the $1$-norm is estimated using the algorithm in [10]. We will use $expm{\_}multiply$ from scipy for $\operatorname{GL}(n)$ and $\mathrm{SO}(n)$, while we will show a customized implementation based on efficient estimates of the $1$-norm of the operator $\mathrm{P}$ for Stiefel and flag manifolds in section 9 could speed up the algorithm significantly, without the higher $p$ estimates.
2.2. Symmetric Bilinear forms
As we will discuss pseudo-Riemannian metric, it is appropriate to review some quick facts about bilinear forms, which could be considered as inner products without the positive-definite requirement. The main results that we need are about subspaces. A symmetric bilinear form, or just form on a vector space $\mathcal{E}$ is a scalar-valued bilinear map from $\mathcal{E}\times\mathcal{E}$ to $\mathbb{R}$, denoted by $\langle v,w\rangle$. For brevity, we will omit the word symmetric as we will only work with symmetric bilinear forms. If we identify $\mathcal{E}$ with $\mathbb{R}^{n}$ and represent the form by a matrix $\mathrm{B}$, then $\langle v,w\rangle=\langle v,w\rangle_{\mathrm{B}}=v^{\mathsf{T}}\mathrm{B}w$ for $v,w\in\mathcal{E}$. It is nondegenerate (or we also say $\mathcal{E}$ is nondegenerate) if $\mathrm{B}$ is nonsingular, or $\langle v,w\rangle=0$ for all $w\in\mathcal{E}$ implies $v=0$. This latter formulation is useful if $\mathcal{E}$ is a matrix space. We can restrict the form $\langle$,$\rangle$ , ⟩ to a subspace $\mathcal{V}\subset\mathcal{E}$. In contrast to the positive-definite case, a form may be degenerate on $\mathcal{V}$ even if it is nondegenate on $\mathcal{E}$. We define the orthogonal complement $\mathcal{V}_{{}_{\perp}}$ of $\mathcal{V}$ as follows
(2) $\mathcal{V}_{{}_{\perp}}=\{w\in\mathcal{E}|\;\langle v,w\rangle=0\text{ for all }v\in\mathcal{V}\}.$
Then $\mathcal{V}_{{}_{\perp}}$ may intersect $\mathcal{V}$ nontrivially. However, we have a generalization of several well-known results for inner product spaces below
Lemma 1.
A subspace $\mathcal{V}\subset\mathcal{E}$ is nondegenerate if and only if $\mathcal{V}_{{}_{\perp}}\cap V=\{0\}$. In that case, we have a unique decomposition $\mathcal{E}=\mathcal{V}\oplus\mathcal{V}_{{}_{\perp}}$, or any vector $w\in\mathcal{E}$ could be written in a unique way as $w=w_{\mathcal{V}}+w_{\mathcal{V}_{\perp}}$, with $w_{\mathcal{V}}\in\mathcal{V},w_{\mathcal{V}_{\perp}}\in\mathcal{V}_{\perp}$. The map $\Pi_{\mathcal{V}}:w\mapsto w_{\mathcal{V}}$ is a projection, the orthogonal projection to $\mathcal{V}$. It satisfies
(3) $\langle\Pi_{\mathcal{V}}w,v\rangle=\langle w,v\rangle.$
Proof.
The equivalence of nondegeneracy of $\mathcal{V}$ and $\mathcal{V}_{{}_{\perp}}\cap V=\{0\}$ is just a rewording of the definitions. As $\mathcal{V}$ is nondegenerate, given a basis $\{v_{i}\}_{i=1}^{d}$, of $\mathcal{V}$, with $d=\dim\mathcal{V}$, let $C=C_{ij}=(\langle v_{i},v_{j}\rangle)_{i,j=1}^{d}$, and $v^{j}=\sum_{i}(C^{-1})_{ji}v_{i}$, then the set $\{v^{j}\}_{j=1}^{d}\subset\mathcal{V}$ satisfies $\langle v_{k},v^{j}\rangle=(C^{-1}C)_{jk}=\delta_{jk}$.
Define $w_{\mathcal{V}}=\sum_{i}\langle v^{i},w\rangle v_{i}\in\mathcal{V}$. Then we can check $\langle w_{\mathcal{V}},v^{j}\rangle=\sum_{i}\langle v^{i},w\rangle\delta_{ij}=\langle v^{j},w\rangle$. By linearity, $\langle w_{\mathcal{V}},v\rangle=\langle w,v\rangle$ for all $v\in\mathcal{V}$, or $\langle w-w_{\mathcal{V}},v\rangle=0$, hence $w-w_{\mathcal{V}}\in\mathcal{V}_{{}_{\perp}}$, and $w=w_{\mathcal{V}}+(w-w_{\mathcal{V}})\in\mathcal{V}\oplus\mathcal{V}_{{}_{\perp}}$. ∎
2.3. Manifolds, geodesics, and parallel transport
To keep the paper self-contained, we will try to minimize the assumed prerequisite from Lie theory and matrix manifolds.
For the background on matrix manifolds, the reader can consult [5, 7, 1]. In the embedded setup, we consider a smooth subset $\mathcal{M}$ of a matrix space $\mathcal{E}=\mathbb{R}^{n\times m}$, a manifold. We equipped $\mathcal{M}$ with a pseudo-Riemannian metric (also called semi-Riemannian metric) $\mathbf{g}$, which is a smooth scalar-valued bilinear form $\langle,\rangle_{X,\mathbf{g}}$ on the tangent space of $\mathcal{M}$ at each $X\in\mathcal{M}$. We write $\langle,\rangle_{\mathbf{g}}$ if the point $X$ is understood. The bilinear form is assumed to be nondegenerate. When it is positive-definite, we have a Riemannian metric, otherwise the form could have negative but nonzero eigenvalues.
A vector field $\mathtt{Y}$ on $\mathcal{M}$ could be considered a $\mathcal{E}=\mathbb{R}^{n\times m}$-valued function, such that for each $X\in\mathcal{M}$, $\mathtt{Y}(X)$ is tangent to $\mathcal{M}$. If $\mathtt{Z}$ is another vector field, we can take the directional derivative $\mathrm{D}_{\mathtt{Y}}\mathtt{Z}$, considering $\mathtt{Z}$ as an $\mathcal{E}$-valued function. However, $\mathrm{D}_{\mathtt{Y}}\mathtt{Z}$ is usually not a vector field (the same reason as why an acceleration on a sphere is not tangent), but we can adjust $\mathrm{D}_{\mathtt{Y}}\mathtt{Z}$ by adding a term of the form $\Gamma(\mathtt{Y},\mathtt{Z})$ so that $\nabla_{\mathtt{Y}}\mathtt{Z}:=\mathrm{D}_{\mathtt{Y}}\mathtt{Z}+\Gamma(\mathtt{Y},\mathtt{Z})$ is a vector field. The term $\Gamma(\mathtt{Y},\mathtt{Z})$ is called a Christoffel function in [5], while $\nabla_{\mathtt{Y}}\mathtt{Z}$ is called a covariant derivative, or a connection. Note that $\Gamma$ is dependent on $X$, and is bilinear in the two tangent vector variables $\xi=\mathtt{Y}(X),\eta=\mathtt{Z}(X)$ for $\xi,\eta\in T_{X}\mathcal{M}$. We write $\Gamma(X;\xi,\eta)$ when we need to show the variable $X$ explicitly. For example, for the sphere, we can check that $\Gamma(X;\xi,\eta)=X\xi^{\mathsf{T}}\eta$ will make $\mathrm{D}_{\mathtt{Y}}\mathtt{Z}+\Gamma(\mathtt{Y},\mathtt{Z})$ a vector field.
There is more than one way to define covariant derivatives. However, there is a unique connection satisfying two additional conditions, called the Levi-Civita connection, which will be the kind of connection that we consider in this article. The two conditions are torsion-free, or $\Gamma(\mathtt{Y},\mathtt{Z})=\Gamma(\mathtt{Z},\mathtt{Y})$ for two vector fields $\mathtt{Y},\mathtt{Z}$, and metric compatible, that means
(4) $\mathrm{D}_{\mathtt{Y}}\langle\mathtt{Z},\mathtt{Z}\rangle_{\mathbf{g}}=2\langle\mathtt{Z},\nabla_{\mathtt{Y}}\mathtt{Z}\rangle_{\mathbf{g}}.$
A geodesic from $X$ with initial velocity $\xi\in T_{X}\mathcal{M}$ is a curve $\gamma$ on the manifold satisfying the equation $\nabla_{\dot{\gamma}(t)}\dot{\gamma}(t)=0$, with the initial value condition $\gamma(0)=X,\dot{\gamma}(0)=\xi$. The map $v\mapsto\gamma(1)$ is the Riemannian exponential map, which is not necessarily the same as the Lie group exponential introduced later.
If $\eta\in T_{X}\mathcal{M}$ is another tangent vector at $X$, then the parallel transport of $\eta$ along the geodesic $\gamma(t)$ from 00 to $t$ is the vector $\mathtt{T}_{0}^{t}(\gamma)\eta=\Delta(t)$, where $\Delta$ is a vector field over the curve $\gamma$, that means for each $0\leq s\leq t$, $\Delta(s)$ is tangent to $\mathcal{M}$ at $\gamma(s)$. Additionally, $\Delta$ must satisfy the transport equation
(5) $\begin{gathered}\dot{\Delta}(s)+\Gamma(\gamma(s);\dot{\gamma}(s),\Delta(s))=0,\\ \Delta(0)=\eta.\end{gathered}$
The above discussion also holds for an open subset of a vector space, and the standard differential geometry approach [23] would identify open subsets of an abstract manifold $\mathcal{M}$ with open subsets of a vector space via local coordinates. We can then define a metric on $\mathcal{M}$ by a compatibility requirement on the coordinates. From here, we can define connections and parallel transport without the embedding requirement.
[…]
for two vector fields $\mathtt{Y},\mathtt{Z}$, with $\mkern 1.5mu\overline{\mkern-1.5mu\mathtt{Y}\mkern-1.5mu}\mkern 1.5mu$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathtt{Z}\mkern-1.5mu}\mkern 1.5mu$ their horizontal lifts, $\nabla$ and $\mkern 1.5mu\overline{\mkern-1.5mu\nabla\mkern-1.5mu}\mkern 1.5mu$ the Levi-Civita connections on $\mathcal{M}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu$, respectively, while $\overline{\left(\nabla_{\mathtt{Y}}\mathtt{Z}\right)}$ is the horizontal lift of $\nabla_{\mathtt{Y}}\mathtt{Z}$, and $\mathrm{H}$ is the projection to the horizontal spaces $\mathcal{H}_{X}$ at each $X\in\mathcal{M}$.
In the next section, we will define matrix groups, a major topic of this paper. It also provides background for quotient manifolds as examples of submersions.
2.4. Matrix Lie groups, Lie algebra, left-invariant metrics, and matrix quotient spaces
A Lie group is a differentiable manifold with a differentiable group structure. We are interested in subgroups of the generalized linear group $\mathcal{E}=\operatorname{GL}(n)$, which is the (multiplication) group of invertible matrices in $\mathbb{R}^{n\times n}$. Thus, for such a group $\mathrm{G}$, $\operatorname{I}_{n}\in\mathrm{G}$ and $X_{1}^{-1}X_{2}\in\mathrm{G}$ if $X_{1},X_{2}\in\mathrm{G}$.
The tangent space $\mathfrak{g}$ of $\mathrm{G}$ at $\operatorname{I}=\operatorname{I}_{n}$ is a Lie subalgebra of $\mathbb{R}^{n\times n}$. Recall a vector subspace $\mathfrak{g}$ of the square matrice space $\mathcal{E}=\mathbb{R}^{n\times n}$ is called a Lie subalgebra if it is closed under the Lie bracket, that means for two matrices $\mathtt{a},\mathtt{b}$ in $\mathfrak{g}$,
(7) $[\mathtt{a},\mathtt{b}]:=\mathtt{a}\mathtt{b}-\mathtt{b}\mathtt{a}\subset\mathfrak{g}.$
The Lie bracket is anticommutative and satisfies Jacobis identity [7, Chapter 3]. The matrix exponential map $\operatorname{expm}$ maps $\mathfrak{g}$ to $\mathrm{G}$. Conversely, given a Lie subalgebra $\mathfrak{g}$ of $\mathbb{R}^{n\times n}$, we can construct a Lie group with Lie algebra $\mathfrak{g}$. In this instant, $\operatorname{expm}$ corresponds to the Lie exponential map of $\operatorname{GL}(n)$, and its subgroups. We have
(8) $\operatorname{expm}(A)X\operatorname{expm}(-A)=\operatorname{expv}(\operatorname{ad}_{A},X),$
where $\operatorname{ad}_{A}:X\mapsto[A,X]$ and $\operatorname{expv}$ is the exponential action in section 2.1 [7, Chapter 3]. From this, $\operatorname{expm}(A)X\operatorname{expm}(-A)\in\mathfrak{g}$ if $A,X\in\mathfrak{g}$. The tangent space of $\mathrm{G}$ at $X\in\mathrm{G}$ could be identified with $X\mathfrak{g}$, or $\xi\in T_{X}\mathrm{G}$ if and only if $X^{-1}\xi\in\mathfrak{g}$. From here, if $\langle$,$\rangle$ , ⟩ is a nondegenerate form on $\mathfrak{g}$, we can define a pseudo-Riemannian metric on $\mathrm{G}$ by defining for two tangent vectors $\xi,\eta$ at $X$
(9) $\langle\xi,\eta\rangle_{X}=\langle X^{-1}\xi,X^{-1}\eta\rangle,$
which gives $\mathrm{G}$ a left-invariant pseudo-Riemannian metric, as $\langle X_{1}\xi,X_{1}\eta\rangle_{X_{1}X}=\langle\xi,\eta\rangle_{X}$ for $X_{1}\in\mathrm{G}$.
In applications here, we will consider right actions, smooth maps ¯×$\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu\times\mathrm{K}\mapsto\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu$, maping $(X,K)\in\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu\times\mathrm{K}$ to $XK$ that are compatible with group multiplication from the right. The actions are assumed proper such that the relationship $X\sim XK$ defines a manifold structure on the quotient $\mathcal{M}/\mathrm{K}$. This is satisfied for the case of compact groups, which we consider here. In this case, we can use the notion of horizontal lift above to compute connections, geodesics, and parallel transports.
2.5. Notations
The generalized linear group $\operatorname{GL}(n)$ is the group of all invertible matrices in $\mathbb{R}^{n\times n}$, the connected subgroup with positive determinant is $\operatorname{GL}^{+}(n)$, the corresponding Lie algebra $\mathfrak{gl}(n)$ could be identified with $\mathbb{R}^{n\times n}$. The special orthogonal group $\mathrm{SO}(n)$ is the group of orthogonal matrices of determinant $1$, its Lie algebra $\mathfrak{so}(n)$ could be identified as a vector space with the space of antisymmetric matrices $\mathrm{Skew}_{n}$. Its orthogonal complement is the space $\mathrm{Sym}_{n}$ of symmetric matrices. The Stiefel, flag and Grassmann manifolds with notations $\mathrm{St}_{n,p},\mathrm{Gr}_{n,p},\operatorname{Flag}(\vec{d})$ will be introduced later. We write $\mathrm{sym}$ for the symmetrize operator $A_{\mathrm{sym}}=\frac{1}{2}(A+A^{\mathsf{T}})$ for $A\in\mathbb{R}^{n\times}$ and $\mathrm{skew}$ for the antisymmetrize operator $A_{\mathrm{skew}}=\frac{1}{2}(A-A^{\mathsf{T}})$.
3. Transposable spaces, trace and Frobenius forms, and a family of tractable metrics
We now turn to bilinear forms on Lie subalgebras and left-invariant metrics on the corresponding groups.
The trace form is the bilinear map $(\mathtt{a},\mathtt{b})\mapsto\operatorname{Tr}\mathtt{a}\mathtt{b}=:\langle\mathtt{a},\mathtt{b}\rangle_{\mathcal{K}}$ (Green the subscript $\mathcal{K}$ is in honor of Killing, who defined the closely related Killing form). Using the commutative property of trace, we have
(10) $\operatorname{Tr}[\mathtt{a},\mathtt{b}]\mathtt{c}=\operatorname{Tr}\mathtt{b}[\mathtt{c},\mathtt{a}],$
or the trace form is the cyclic invariant. As any strictly upper-triangular matrix $\mathtt{a}$ has $\operatorname{Tr}\mathtt{a}^{2}=0$, the trace form is not positive definite. A symmetric matrix $\mathtt{a}$ has $\operatorname{Tr}\mathtt{a}^{2}\geq 0$, while an antisymmetric matrix has $\operatorname{Tr}\mathtt{a}^{2}\leq 0$. However, the trace form has a very nice property that if it is nondegenerate on $\mathfrak{g}$, it induces a bi-invariant pseudo-Riemannian metric on the corresponding group with closed-form geodesic.
The Frobenius form, or Frobenius inner product $(\mathtt{a},\mathtt{b})\mapsto\langle\mathtt{a},\mathtt{b}\rangle_{\mathcal{F}}=:\operatorname{Tr}\mathtt{a}\mathtt{b}^{\mathsf{T}}$ is positive-definite. However, it may not induce a metric with closed-form geodesics. We will focus on subspaces that transpose to themselves, where we can use the nice properties of trace and Frobenius forms to construct tractable metrics.
Definition 1.
A matrix vector space $\mathcal{V}\subset\mathbb{R}^{n\times n}$ is called transposable if
(11) $\mathcal{V}^{\mathsf{T}}=\{\mathtt{a}^{\mathsf{T}}|\mathtt{a}\in\mathcal{V}\}=\mathcal{V}.$
A transposable Lie subalgebra is a transposable vector space that is a Lie subalgebra.
The spaces of symmetric, antisymmetric, diagonal matrices, and block-diagonal matrices with fixed block sizes are all transposable. If $\mathcal{V}$ is transposable then its symmetric $\mathcal{V}_{\mathrm{sym}}$ and antisymmetric $\mathcal{V}_{\mathrm{skew}}$ parts are subspaces of $\mathcal{V}$.
Transposable Lie subalgebras have an important property. Let us explain it through an example of a space of block matrices. For $k<n$, let $\mathfrak{a}$ be a copy of $\mathbb{R}^{k\times k}$, considered as a subspace of $\mathcal{E}=\mathbb{R}^{n\times n}$ by identifying $\mathfrak{a}$ with the space of upper-left $k\times k$ block diagonal matrices, and $\mathfrak{a}_{{}_{\perp}}$ its Frobenius orthogonal complements. Thus $\mathfrak{a}$ consists of matrice of the form $\begin{bmatrix}A&0\\ 0&0\end{bmatrix}$ with $A\in\mathbb{R}^{k\times k}$, while $\mathfrak{a}_{{}_{\perp}}$ consists of those of the form $\begin{bmatrix}0&B\\ C&D\end{bmatrix}$, with $D\in\mathbb{R}^{(n-k)\times(n-k)},B\in\mathbb{R}^{k\times(n-k)},C\in\mathbb{R}^{(n-k)\times k}$. Notice that $\mathfrak{a}$ and $\mathfrak{d}=\{\begin{bmatrix}0&0\\ 0&D\end{bmatrix}|D\in\mathbb{R}^{(n-k)\times(n-k)}\}$ commutes, while the Lie bracket of $\mathfrak{a}$ or $\mathfrak{d}$ with a matrix of the form $\begin{bmatrix}0&B\\ C&0\end{bmatrix}$ is another matrix of that form. All three spaces $\mathfrak{a},\mathfrak{d}$ and the remaining part are transposable. The splitting of $\mathbb{R}^{n\times n}$ to these three parts is a special case of proposition 1 below where $\mathfrak{g}$ is $\mathbb{R}^{n\times n}$, $\mathfrak{a}$ is the space of top $k\times k$ block matrices and $\mathfrak{d}$ is $\mathfrak{a}_{{}_{\top}}$, the space of bottom $(n-k)\times(n-k)$ matrices. The remaining space with zero diagonal blocks is $\mathfrak{a}_{{}_{\Join}}$. We can add $\mathfrak{a}_{{}_{\Join}}$ to either $\mathfrak{a}$ or $\mathfrak{a}_{{}_{\top}}$ to form the orthogonal complement of the remaining algebra.
Proposition 1.
Assume $\mathfrak{g}$ is transposable. Then the trace form is nondegenerate on $\mathfrak{g}$. Let $\mathfrak{v}\subset\mathfrak{g}$ be a transposable subspace. Let $\mathfrak{v}_{{}_{\perp}}$ be the orthogonal complement of $\mathfrak{v}$ under the trace form $\mathcal{K}$
(12) $\mathfrak{v}_{{}_{\perp}}=\{\mathtt{h}\in\mathfrak{g}|\;\operatorname{Tr}\mathtt{h}\mathfrak{v}=\{0\}\}.$
Then $\mathfrak{v}_{{}_{\perp}}$ is also transposable and is the orthogonal complement under the Frobenius form $\mathcal{F}$, and vice-versa. The projection $\mathtt{p}_{\mathfrak{v}}$ of $\mathfrak{g}$ onto $\mathfrak{v}$ under $\mathcal{K}$ and $\mathcal{F}$ is identical and commutes with transpose, $\mathtt{p}_{\mathfrak{v}}(\mathtt{h}^{\mathsf{T}})=\mathtt{p}_{\mathfrak{v}}(\mathtt{h})^{\mathsf{T}}$ for $\mathtt{h}\in\mathfrak{g}$.
If $\mathfrak{a}$ is a transposable subalgebra, then we have
[…]
and $\mathfrak{a}_{{}_{\Join}}$ is also transposable. The space $\mathfrak{a}_{{}_{\top}}$ defined by either one of the equations below is a transposable Lie subalgebra of $\mathfrak{g}$
(14) $\mathfrak{a}_{{}_{\top}}=\{\mathtt{c}\in\mathfrak{a}_{{}_{\perp}}|\operatorname{Tr}\mathtt{c}[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]=\{0\}\},$
(15) $\mathfrak{a}_{{}_{\top}}=\{\mathtt{c}\in\mathfrak{a}_{{}_{\perp}}|[\mathtt{c},\mathfrak{a}]=\{0\}.$
The correspondence $\mathfrak{a}\mapsto\mathfrak{a}_{{}_{\top}}$ is involutive, $(\mathfrak{a}_{{}_{\top}})_{{}_{\top}}=\mathfrak{a}$. We have $[\mathfrak{a},\mathfrak{a}_{{}_{\Join}}]\subset\mathfrak{a}_{{}_{\Join}}$ and $[\mathfrak{a}_{{}_{\top}},\mathfrak{a}_{{}_{\Join}}]\subset\mathfrak{a}_{{}_{\Join}}$ and we have an orthogonal decomposition under either $\mathcal{K}$ or $\mathcal{F}$
(16) $\mathfrak{g}=\mathfrak{a}\oplus\mathfrak{a}_{{}_{\Join}}\oplus\mathfrak{a}_{{}_{\top}}.$
Proof.
Note for $\mathtt{g}_{1},\mathtt{g}_{2}\in\mathfrak{g}$, $\langle\mathtt{g}_{1},\mathtt{g}_{2}\rangle_{\mathcal{F}}=\langle\mathtt{g}_{1}^{\mathsf{T}},\mathtt{g}_{2}\rangle_{\mathcal{K}}$, hence, if $\langle\mathtt{g}_{1},\mathtt{g}_{2}\rangle_{\mathcal{K}}=0$ for all $\mathtt{g}_{2}\in\mathfrak{g}$, then $\langle\mathtt{g}_{1}^{\mathsf{T}},\mathtt{g}_{2}\rangle_{\mathcal{F}}=0$ for all $\mathtt{g}_{2}$, therefore $\mathtt{g}_{1}^{\mathsf{T}}=0$, or $\mathtt{g}_{1}=0$. Thus, the trace form is nondegenerate on $\mathfrak{g}$.
The next statement is a consequence of $\mathfrak{v}=\mathfrak{v}^{\mathsf{T}}$, which implies $\mathfrak{v}_{{}_{\perp}}=\{\mathtt{h}\in\mathfrak{g}|\;\operatorname{Tr}\mathtt{h}\mathfrak{v}^{\mathsf{T}}=\{0\}\}$, or $\mathfrak{v}_{{}_{\perp}}$ is also the orthogonal complement under the Frobenius inner product. Using $\operatorname{Tr}a^{\mathsf{T}}b=\operatorname{Tr}ab^{\mathsf{T}}$ for two matrices $a,b$, we get $\mathfrak{v}_{{}_{\perp}}$ is transposable. From here, the decomposition $\mathtt{g}=\mathtt{g}_{\mathfrak{v}}+\mathtt{g}_{\mathfrak{v}_{{}_{\perp}}}$ under the trace form is also the decomposition under the Frobenius form. Thus, the projection under one form is the projection under the other. Transposing the decomposition to $\mathtt{g}^{\mathsf{T}}=\mathtt{g}_{\mathfrak{v}}^{\mathsf{T}}+\mathtt{g}_{\mathfrak{v}_{{}_{\perp}}}^{\mathsf{T}}$ then use the uniqueness of this decompostion, we get $(\mathtt{g}^{\mathsf{T}})_{\mathfrak{v}}=(\mathtt{g}_{\mathfrak{v}})^{\mathsf{T}}$.
To show $[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]\subset\mathfrak{a}_{{}_{\perp}}$, let $\mathtt{c}\in\mathfrak{a}$, using eq. 10, we compute $\operatorname{Tr}\mathtt{c}[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]=\operatorname{Tr}\mathfrak{a}_{{}_{\perp}}[\mathtt{c},\mathfrak{a}]=0$ as $[\mathtt{c},\mathfrak{a}]\subset\mathfrak{a}$ since $\mathfrak{a}$ is a subalgebra. For $\mathtt{a}\in\mathfrak{a},\mathtt{b}\in\mathfrak{a}_{{}_{\perp}}$, $[\mathtt{a},\mathtt{b}]^{\mathsf{T}}=-[\mathtt{a}^{\mathsf{T}},\mathtt{b}^{\mathsf{T}}]\subset[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]$, proving $\mathfrak{a}_{{}_{\Join}}^{\mathsf{T}}=\mathfrak{a}_{{}_{\Join}}$. We have eq. 15 implies eq. 14 because $\operatorname{Tr}\mathtt{c}[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]=\operatorname{Tr}\mathfrak{a}_{{}_{\perp}}[\mathtt{c},\mathfrak{a}]$. For the reverse, since $\mathtt{c}\in\mathfrak{a}_{{}_{\perp}}$, $[\mathtt{c},\mathfrak{a}]\in\mathfrak{a}_{{}_{\perp}}$ from eq. 13. As $\mathfrak{a}_{{}_{\perp}}$ is transposable, the trace form is nondegenerate on $\mathfrak{a}_{{}_{\perp}}$, and this shows $[\mathtt{c},\mathfrak{a}]=0$.
Finally, using eq. 15 and $(\mathfrak{a}_{{}_{\top}})_{{}_{\perp}}=\mathfrak{a}\oplus\mathfrak{a}_{{}_{\Join}}$, we get $\mathfrak{a}\mapsto\mathfrak{a}_{{}_{\top}}$ is involutive. ∎
From now on, when $\mathfrak{v}$ is transposable, we will use the notation $\mathtt{p}_{\mathfrak{v}}$ to denote the projection under either the Frobenius or trace form, and write $w_{\mathfrak{v}}$ for $\mathtt{p}_{\mathfrak{v}}(w)$ for $w\in\mathcal{E}$. We denote by $\mathfrak{v}_{{}_{\perp}}$ the orthogonal complement from either the trace or Frobenius form.
We now describe projections to several transposable subspaces that we will encounter. A projection to a block diagonal subspace is just setting the entries outside of those blocks to zero. A projection to the space of symmetric matrices is just by applying the symmetrize operator, $A\mapsto A_{\mathrm{sym}}$, projection to the space of antisymmetric matrices is by applying the antisymmetrize operator.
Let $\mathrm{G}\subset\mathcal{E}:=\mathbb{R}^{n\times n}$ be a connected matrix Lie group with Lie algebra $\mathfrak{g}$.
Theorem 1.
Let $\mathrm{G}\subset\mathbb{R}^{n\times n}$ be a connected matrix Lie group with a transposable Lie subalgebra $\mathfrak{g}$, and $\mathfrak{a}\subset\mathfrak{g}$ is a transposable Lie subalgebra. Let $\vec{\beta}=(\beta_{0},\beta_{1})\in\mathbb{R}^{2}$ where $\beta_{0}\beta_{1}\neq 0$, and set $\beta=\frac{\beta_{1}}{\beta_{0}}$. Define the symmetric bilinear form on $\mathfrak{g}$
(17) $\langle\mathtt{g},\mathtt{h}\rangle_{\vec{\beta}}=\beta_{0}(\operatorname{Tr}(\mathtt{h}\mathtt{g})-\operatorname{Tr}(\mathtt{h}_{\mathfrak{a}}\mathtt{g}_{\mathfrak{a}}))-\beta_{1}\operatorname{Tr}(\mathtt{h}_{\mathfrak{a}}\mathtt{g}_{\mathfrak{a}}).$
The form is nondegenerate, which defines a left-invariant pseudo-Riemannian metric on the Lie group $\mathrm{G}$ by eq. 9. For two vector fields $\mathtt{Y},\mathtt{Z}$, the Levi-Civita connection of this metric is given by $\nabla_{\mathtt{Y}}\mathtt{Z}=\mathrm{D}_{\mathtt{Y}}\mathtt{Z}+\Gamma(\mathtt{Y},\mathtt{Z})$ where the Christoffel function $\Gamma$ for two tangent vectors $\xi,\eta\in T_{X}\mathrm{G}$ is given by
(18) $\Gamma(\xi,\eta)=-\frac{1}{2}(\xi X^{-1}\eta+\eta X^{-1}\xi)+\frac{1+\beta}{2}X\left([(X^{-1}\xi)_{\mathfrak{a}},X^{-1}\eta]+[(X^{-1}\eta)_{\mathfrak{a}},X^{-1}\xi]\right).$
Further, the last term above is in $X\mathfrak{a}_{{}_{\Join}}$, as for $\mathtt{a},\mathtt{b}\in\mathfrak{g}$, we have
(19) $[\mathtt{a}_{\mathfrak{a}},\mathtt{b}]+[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]\in\mathfrak{a}_{{}_{\Join}}.$
The geodesic $\gamma(t)$ starting at $\gamma(0)=X\in\mathrm{G}$ with initial velocity $\dot{\gamma}(0)=\xi$ is given by
(20) $\gamma(t)=X\operatorname{expm}(t\left(-(1+\beta)(X^{-1}\xi)_{\mathfrak{a}}+X^{-1}\xi\right))\operatorname{expm}(t(1+\beta)(X^{-1}\xi)_{\mathfrak{a}}).$
Set $\mathtt{a}=X^{-1}\xi$, let $\eta\in T_{X}\mathcal{M}$ be another tangent vector, then the parallel transport of $\eta$ along the geodesics $\gamma$ to $X_{t}=\gamma(t)$ with initial point and velocity $(X,\xi)$ is
(21) $\begin{gathered}\mathtt{T}(\gamma)_{0}^{t}\eta=X\operatorname{expm}(t\left(-(1+\beta)\mathtt{a}_{\mathfrak{a}}+\mathtt{a}\right))\operatorname{expv}(t\mathrm{P}_{\mathtt{a}},X^{-1}\eta)\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}})\\ \text{ where for }\mathtt{b}\in\mathfrak{g},\mathrm{P}_{\mathtt{a}}:\mathtt{b}\mapsto\frac{1}{2}\left([\mathtt{b},\mathtt{a}]+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}])\right).\end{gathered}$
In the Frobenius inner product, the adjoint of the operator $\mathrm{P}_{\mathtt{a}}$ is given by
(22) $\mathrm{P}_{\mathtt{a}}^{\mathsf{T}}\mathtt{b}=\frac{1}{2}\left([\mathtt{b},\mathtt{a}^{\mathsf{T}}]+(1+\beta)([\mathtt{a}_{\mathfrak{a}}^{\mathsf{T}},\mathtt{b}]-[\mathtt{b},\mathtt{a}^{\mathsf{T}}]_{\mathfrak{a}})\right).$
The operator $\mathrm{P}_{\mathtt{a}}$ is antisymmetric in the bilinear pairing eq. 17. For $\mathtt{a},\mathtt{b},\mathtt{c}\in\mathfrak{g}$
(23) $\langle\mathrm{P}_{\mathtt{a}}\mathtt{b},\mathtt{c}\rangle_{\vec{\beta}}+\langle\mathrm{P}_{\mathtt{a}}\mathtt{c},\mathtt{b}\rangle_{\vec{\beta}}=0.$
It is known $\mathtt{T}(\gamma)_{0}^{t}$ is an isometry, which also follows from eq. 23.
Proof.
Set $\mathcal{I}_{\vec{\beta}}:\omega\mapsto\beta_{0}(\omega-\omega_{\mathfrak{a}})^{\mathsf{T}}-\beta_{1}\omega_{\mathfrak{a}}^{\mathsf{T}}$, then $\langle\omega_{1},\omega_{2}\rangle_{\vec{\beta}}=\langle\omega_{1},\mathcal{I}_{\vec{\beta}}\omega_{2}\rangle_{\mathcal{F}}$. As an operator on $\mathcal{E}=\mathbb{R}^{n\times n}$, the operator $\mathcal{I}_{\vec{\beta}}\circ\mathsf{T}$ has two eigenvalues, $\beta_{0}$ with eigenspace $(\operatorname{I}_{\mathcal{E}}-\mathrm{p}_{\mathfrak{a}})\mathcal{E}$, (recall $\mathrm{p}_{\mathfrak{a}}$ is the projection to $\mathfrak{a}$) and $-\beta_{1}$ with eigenspace $\mathfrak{a}$, both are nonzero by assumption. Thus, the operator $\mathcal{I}_{\vec{\beta}}$ and the form $\langle\rangle_{\vec{\beta}}$ are nondegenerate.
For eq. 19, decompose $\mathtt{a}=\mathtt{a}_{\mathfrak{a}}+\mathtt{a}_{\mathfrak{a}_{{}_{\perp}}},\mathtt{b}=\mathtt{b}_{\mathfrak{a}}+\mathtt{b}_{\mathfrak{a}_{{}_{\perp}}}$, and substitute in, note $[\mathtt{a}_{\mathfrak{a}},\mathtt{b}_{\mathfrak{a}}]+[\mathtt{b}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]=0$, the remaining part is in $\mathfrak{a}_{{}_{\Join}}$ by eq. 13.
To show eq. 18 defines a connection, we need to show $X^{-1}(\nabla_{\mathtt{Y}}\mathtt{Z})_{X}\in\mathfrak{g}$, and we only need to show the following expression (dropping the variable X in the functions for brevity) is in $\mathfrak{g}$ when evaluated at $X\in\mathrm{G}$, as the remaining terms are already in $\mathfrak{g}$
$X^{-1}\mathrm{D}_{\mathtt{Y}}\mathtt{Z}-\frac{1}{2}(X^{-1}\mathtt{Y}X^{-1}\mathtt{Z}+X^{-1}\mathtt{Z}X^{-1}\mathtt{Y})\in\mathfrak{g}.$
Here, set $\mathtt{y}(X)=X^{-1}\mathtt{Y}(X),\mathtt{z}(X)=X^{-1}\mathtt{Z}(X)$, then $\mathtt{y}$ and $\mathtt{z}$ are two $\mathfrak{g}$-valued functions on $\mathrm{G}$, the expression becomes
(24) $X^{-1}(X\mathtt{y}\mathtt{z}+X\mathrm{D}_{\mathtt{Y}}\mathtt{z}-\frac{1}{2}(\mathtt{y}\mathtt{z}+\mathtt{z}\mathtt{y}))=\mathrm{D}_{\mathtt{Y}}\mathtt{z}+\frac{1}{2}[\mathtt{y},\mathtt{z}]\in\mathfrak{g}$
where $\mathrm{D}_{\mathtt{Y}}\mathtt{z}\in\mathfrak{g}$ because of linearity, while $[\mathtt{y},\mathtt{z}]\in\mathfrak{g}$ because $\mathfrak{g}$ is closed under the Lie bracket. As $\Gamma$ is symmetric in $\mathtt{Y}$ and $\mathtt{Z}$, it remains to check for metric compatibility, or to show $C:=\mathrm{D}_{\mathtt{Y}}\langle\mathtt{Z},\mathtt{Z}\rangle_{\vec{\beta}}-2\langle\mathtt{Z},\nabla_{\mathtt{Y}}\mathtt{Z}\rangle_{\vec{\beta}}$ vanishes assuming eq. 18. Expand, we have
$\begin{gathered}C=\beta_{0}\mathrm{D}_{\mathtt{Y}}(\operatorname{Tr}\mathtt{z}^{2})-(\beta_{0}+\beta_{1})\mathrm{D}_{\mathtt{Y}}(\operatorname{Tr}\mathtt{z}_{\mathfrak{a}}^{2})\\ -2\beta_{0}\operatorname{Tr}\mathtt{z}X^{-1}\nabla_{\mathtt{Y}}\mathtt{Z}+2(\beta_{0}+\beta_{1})\operatorname{Tr}\mathtt{z}_{\mathfrak{a}}(X^{-1}\nabla_{\mathtt{Y}}\mathtt{Z})_{\mathfrak{a}}.\end{gathered}$
Using eq. 24 to simplify $X^{-1}\nabla_{\mathtt{Y}}\mathtt{Z}$, the second line becomes
$\begin{gathered}-\beta_{0}\left(2\operatorname{Tr}\mathtt{z}\mathrm{D}_{\mathtt{Y}}\mathtt{z}+\operatorname{Tr}\mathtt{z}[\mathtt{y},\mathtt{z}]+(1+\beta)\operatorname{Tr}\mathtt{z}\left([\mathtt{y}_{\mathfrak{a}},\mathtt{z}]+[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]\right)\right)\\ +(\beta_{0}+\beta_{1})\left(2\operatorname{Tr}\mathtt{z}_{\mathtt{a}}\mathrm{D}_{\mathtt{Y}}\mathtt{z}_{\mathtt{a}}+\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]_{\mathtt{a}}+(1+\beta)\operatorname{Tr}\mathtt{z}_{\mathtt{a}}\left([\mathtt{y}_{\mathfrak{a}},\mathtt{z}]+[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]\right)\right)_{\mathtt{a}}.\end{gathered}$
The last term vanishes by eq. 19. Putting together and cancelling terms with $\mathrm{D}_{\mathtt{Y}}\mathtt{z}$
$\begin{gathered}C=-\beta_{0}\left(\operatorname{Tr}\mathtt{z}[\mathtt{y},\mathtt{z}]+(1+\beta)\operatorname{Tr}\mathtt{z}\left([\mathtt{y}_{\mathfrak{a}},\mathtt{z}]+[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]\right)\right)\\ +(\beta_{0}+\beta_{1})\left(\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]_{\mathtt{a}}\right)\\ =-(\beta_{0}+\beta_{1})\operatorname{Tr}\mathtt{z}[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]+(\beta_{0}+\beta_{1})\left(\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]_{\mathtt{a}}\right)\\ =-(\beta_{0}+\beta_{1})\left(\operatorname{Tr}\mathtt{z}[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]-\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]\right)=0,\end{gathered}$
where we note $\operatorname{Tr}\mathtt{z}[\mathtt{y},\mathtt{z}]=0=\operatorname{Tr}\mathtt{z}[\mathtt{y}_{\mathfrak{a}},\mathtt{z}]$ by eq. 10, $\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]=\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]_{\mathfrak{a}}$ since $\mathrm{p}_{\mathfrak{a}}$ is a projection, then finish the cancelation by eq. 10 again.
To verify that eq. 20 satisfies the geodesic equation, with $\mathtt{a}=X^{-1}\xi$, we have
(25) $\dot{\gamma}(t)=X\operatorname{expm}(t\left(-(1+\beta)\mathtt{a}_{\mathfrak{a}}+\mathtt{a}\right))\mathtt{a}\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}}),$
(26) $\ddot{\gamma}(t)=X\operatorname{expm}(t\left(-(1+\beta)\mathtt{a}_{\mathfrak{a}}+\mathtt{a}\right))\left(\mathtt{a}^{2}+(1+\beta)[\mathtt{a},\mathtt{a}_{\mathfrak{a}}]\right)\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}}).$
From here $\gamma^{-1}\dot{\gamma}=U^{-1}\mathtt{a}U$ for $U=\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}})$, while
$D:=\gamma(t)^{-1}\Gamma(\dot{\gamma}(t),\dot{\gamma}(t))_{X=\gamma(t)}=-U^{-1}\mathtt{a}^{2}U+(1+\beta)U^{-1}[\mathtt{a}_{\mathfrak{a}},\mathtt{a}]U)$
since from the expression for $U$, $U^{-1}\mathtt{a}_{\mathfrak{a}}U=\mathtt{a}_{\mathfrak{a}}$, thus, $D$ simplies to the above. This shows $\ddot{\gamma}(t)+\Gamma(\dot{\gamma},\dot{\gamma})=0$, hence $\gamma(t)$ is the geodesic with $\gamma(0)=X,\dot{\gamma}(0)=X\mathtt{a}=\xi$.
For parallel transport, assume the vector field $\Delta(t)$ over the curve $\gamma(t)$ satisfies the transport equation $\dot{\Delta}+\Gamma(\dot{\gamma},\Delta)=0.$ Let $W(t)=U(t)\gamma(t)^{-1}\Delta(t)U(t)^{-1}$, where $U(t)=\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}})$. We will show $W(t)$ satisfies a linear differential equation with constant coefficients. Express $\Delta$ in terms of $W(t)$ and take time derivative
$\begin{gathered}\Delta=\gamma U^{-1}WU,\\ \dot{\Delta}=\dot{\gamma}U^{-1}WU+(1+\beta)\gamma U^{-1}[W,\mathtt{a}_{\mathfrak{a}}]U+\gamma U^{-1}\dot{W}U\\ =\gamma U^{-1}(\mathtt{a}W+(1+\beta)[W,\mathtt{a}_{\mathfrak{a}}]+\dot{W})U,\end{gathered}$
since $\dot{\gamma}=\gamma U^{-1}\mathtt{a}U$, thus, $\dot{\gamma}U^{-1}WU=\gamma U^{-1}\mathtt{a}UU^{-1}WU=\gamma U^{-1}\mathtt{a}WU$. Next,
$\Gamma(\dot{\gamma},\Delta)=-\frac{1}{2}\gamma U^{-1}(\mathtt{a}W+W\mathtt{a})U+\frac{1+\beta}{2}\gamma U^{-1}\left([\mathtt{a}_{\mathfrak{a}},W]+[W_{\mathfrak{a}},\mathtt{a}]\right)U.$
From here, $U\gamma^{-1}(\dot{\Delta}+\Gamma(\dot{\gamma},\Delta))U^{-1}=0$ gives the equation for $W$
$\dot{W}+\frac{1}{2}[\mathtt{a},W]+\frac{1+\beta}{2}\left([W_{\mathfrak{a}},\mathtt{a}]+[W,\mathtt{a}_{\mathfrak{a}}]\right)=0.$
Thus, $W(t)=\operatorname{expv}(t\mathrm{P}_{\mathtt{a}},W_{0})$ with $W_{0}=X^{-1}\eta$, implying eq. 21.
To prove eq. 22, we use the equality $\operatorname{Tr}\mathtt{b}_{1}^{\mathsf{T}}[\mathtt{b}_{2},\mathtt{g}]=\operatorname{Tr}\mathtt{b}_{2}^{\mathsf{T}}[\mathtt{b}_{1},\mathtt{g}^{\mathsf{T}}]$ for $\mathtt{b}_{1},\mathtt{b}_{2},\mathtt{g}\in\mathbb{R}^{n\times n}$ by taking transpose then use eq. 10. Thus, for $\mathtt{c}\in\mathcal{E}=\mathbb{R}^{n\times n}$,
$\begin{gathered}\operatorname{Tr}\mathtt{c}^{\mathsf{T}}\left(\frac{1}{2}[\mathtt{b},\mathtt{a}]+\frac{1+\beta}{2}\left(-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]-[\mathtt{b},\mathtt{a}_{\mathfrak{a}}]\right)\right)\\ =\frac{1}{2}\operatorname{Tr}\mathtt{b}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]-\frac{1+\beta}{2}\left(\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]+\operatorname{Tr}\mathtt{b}^{\mathsf{T}}[\mathtt{c},\mathtt{a}_{\mathfrak{a}}^{\mathsf{T}}]\right).\end{gathered}$
Using $\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]=\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]_{\mathfrak{a}}=\operatorname{Tr}\mathtt{b}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]_{\mathfrak{a}}$, we get the expression for $\mathrm{P}^{\mathsf{T}}_{\mathtt{a}}$.
We prove eq. 23 by a direct calculation, note $\operatorname{Tr}\mathtt{g}_{\mathfrak{a}}\mathtt{c}_{\mathfrak{a}}=\operatorname{Tr}\mathtt{g}\mathtt{c}_{\mathfrak{a}}$ for $\mathtt{g}\in\mathfrak{g}$
$\begin{gathered}\frac{1}{\beta_{0}}\langle\mathrm{P}_{\mathtt{a}}\mathtt{b},\mathtt{c}\rangle_{\vec{\beta}}=\operatorname{Tr}\frac{1}{2}\left([\mathtt{b},\mathtt{a}]\mathtt{c}+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]\mathtt{c}-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]\mathtt{c})\right)\\ -(1+\beta)\operatorname{Tr}\left(\frac{1}{2}\left([\mathtt{b},\mathtt{a}]\mathtt{c}_{\mathfrak{a}}+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]\mathtt{c}_{\mathfrak{a}}-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]\mathtt{c}_{\mathfrak{a}})\right)\right)\\ =\operatorname{Tr}\frac{1}{2}\left(\mathtt{b}[\mathtt{a},\mathtt{c}]+(1+\beta)(\mathtt{b}[\mathtt{c},\mathtt{a}_{\mathfrak{a}}]-\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}])\right)\\ -(1+\beta)\operatorname{Tr}\left(\frac{1}{2}\left(\mathtt{b}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}]+(1+\beta)(\mathtt{b}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]-\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}])\right)\right)\\ =\frac{1}{2}\operatorname{Tr}\mathtt{b}[\mathtt{a},\mathtt{c}]+\frac{1+\beta}{2}\operatorname{Tr}(\mathtt{b}[\mathtt{c},\mathtt{a}_{\mathfrak{a}}]-\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}]-\mathtt{b}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}])\\ -\frac{(1+\beta)^{2}}{2}\operatorname{Tr}(\mathtt{b}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]-\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}]).\end{gathered}$
It remains to prove the following, as by eq. 10, all terms on the next to last line behave antisymmetrically when we change the positions of $\mathtt{b}$ and $\mathtt{c}$
$\operatorname{Tr}\mathtt{b}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]-\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}]+\operatorname{Tr}\mathtt{c}[\mathtt{b}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]-\operatorname{Tr}\mathtt{c}_{\mathfrak{a}}[\mathtt{a},\mathtt{b}_{\mathfrak{a}}]=0.$
The first and last term cancel, since $\operatorname{Tr}\mathtt{b}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]=\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]$ as $\mathfrak{a}$ is a subalgebra, while $\operatorname{Tr}\mathtt{c}_{\mathfrak{a}}[\mathtt{a},\mathtt{b}_{\mathfrak{a}}]=\operatorname{Tr}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}]\mathtt{b}_{\mathfrak{a}}=\operatorname{Tr}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]\mathtt{b}_{\mathfrak{a}}$ as $[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}_{{}_{\perp}}}]\subset\mathfrak{a}_{{}_{\perp}}$. The middle two terms cancel similarly, switching $\mathtt{b}$ and $\mathtt{c}$. ∎
3.1. Riemannian metrics from the family
The case $\beta=-1$ is particularly simple, as we see the metric eq. 17 reduces to a multiple of the trace form $\beta_{0}\operatorname{Tr}\mathtt{h}\mathtt{g}$, the connection $\Gamma(\xi,\eta)$ reduces to the first two terms, the geodesic is $X\operatorname{expm}(tX^{-1}\xi)$, or just $\operatorname{expm}(t\xi)$ if $X=\operatorname{I}_{\mathcal{E}}$, the operator $\mathrm{P}_{\mathtt{a}}$ is just $-\frac{1}{2}\operatorname{ad}_{\mathtt{a}}$, and using eq. 8, the parallel transport is $X\operatorname{expm}(\frac{t}{2}\mathtt{a})X^{-1}\eta\operatorname{expm}(\frac{t}{2}\mathtt{a})$. For compact groups, this is known.
However, the trace form metric may not be Riemannian. The family of metrics in theorem 1 could be considered a deformation of the trace form using the Lie algebra $\mathfrak{a}$. For certain choices of the subgroup $\mathfrak{a}$ and parameters $\vec{\beta}$, the metric becomes Riemannian. The condition for this could be analyzed by studying the eigenvalues of $\mathcal{I}_{\vec{\beta}}$ above, which could be written as
(27) $\mathcal{I}_{\vec{\beta}}:\mathsf{T}\circ\left(\beta_{0}(\rm{I}_{\mathcal{E}}-\mathrm{p}_{\mathfrak{a}})-\beta_{1}\circ\mathrm{p}_{\mathfrak{a}}\right)=\left(\beta_{0}(\rm{I}_{\mathcal{E}}-\mathrm{p}_{\mathfrak{a}})-\beta_{1}\circ\mathrm{p}_{\mathfrak{a}}\right)\circ\mathsf{T}.$
Note that the transpose $\mathsf{T}$ has eigenvalues ±$\pm 1$ on the spaces of symmetric and antisymmetric matrices. Write $\mathcal{V}_{\mathrm{sym}}$ and $\mathcal{V}_{\mathrm{skew}}$ for the symmetric and antisymmetric components of a transposable subspace $\mathcal{V}\subset\mathcal{E}$, the possible eigenpairs of $\mathcal{I}_{\vec{\beta}}$ are $(\beta_{0},(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}),(-\beta_{0},(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}),(-\beta_{1},\mathfrak{a}_{\mathrm{sym}}),(\beta_{1},(\mathfrak{a})_{\mathrm{skew}})$. The Riemannian condition implies at least one of the first two, and at least one of the last two eigenspaces vanish, assuming the appropriate signs of $\beta_{0}$ and $\beta_{1}$. These are the possible cases
  1. (1)
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}=\{0\},(\mathfrak{a})_{\mathrm{sym}}=\{0\}$. In this case $\mathfrak{g}_{\mathrm{sym}}=\{0\}$, or $\mathrm{G}$ is a compact Lie subgroup of $\mathrm{SO}(n)$. In this case $\beta_{0}<0,\beta_{1}>0$. An example is $\mathrm{SO}(n)$ itself in section 5.
  2. (2)
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}=\{0\},(\mathfrak{a})_{\mathrm{skew}}=\{0\}$. For example, $\mathrm{G}$ could be a product of of a compact subgroup (for $\mathfrak{a}_{{}_{\perp}}$) and a diagonal group (for $\mathfrak{a}$). We need $\beta_{0}<0,\beta_{1}>0$.
  3. (3)
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}=\{0\},(\mathfrak{a})_{\mathrm{sym}}=\{0\}$. This shows $\mathfrak{a}=\mathfrak{g}_{\mathrm{skew}}$. An example is $\operatorname{GL}(n)$ in [28, 18], reviewwed under our point of view with parallel transport in section 4. The case of $\mathrm{Sp}(n)$ in [6] could be treated similarly. In this case $\beta_{0}>0,\beta_{1}>0$. Except for the case $\mathfrak{g}_{\mathrm{sym}}=\{0\}$ already considered, the trace form is nondefinite, it is positive on $\mathfrak{g}_{\mathrm{sym}}$ and negative on $\mathfrak{g}_{\mathrm{skew}}$. The metric in eq. 17 modify the trace form on the antisymmetric part by a negative multiplier $-\beta_{1}$. This is also known in [3].
  4. (4)
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}=\{0\},(\mathfrak{a})_{\mathrm{skew}}=\{0\}$. In this case, $\mathfrak{g}\subset\mathcal{E}_{\mathrm{sym}}$, an example is a group of commutative symmetric matrices.
A generalization of theorem 1 is to replace $\beta_{1}$ with a symmetric operator $\mathrm{B}_{1}$ on $\mathfrak{a}$, such that $[\mathrm{B}_{1}\mathtt{a},\mathtt{b}]=\mathrm{B}_{1}[\mathtt{a},\mathtt{b}]$ for $\mathtt{a},\mathtt{b}\in\mathfrak{a}$ ($\mathrm{B}_{1}$ is called an intertwining operator), such that $\mathrm{B}_{1}$ has nonzero eigenvalue, we can check all the arguments in the proof of theorem 1 still hold. For $\mathfrak{so}(n)$ (a simple Lie algbra), such an operator must be a scalar multiplication. If $\mathfrak{a}$ is composed of several simple components, $\mathrm{B}_{1}$ could be constructed by applying different multipliers on different components. We discuss this briefly in section 4.
4. The Generalized Linear group with positive determinant
In this case, the group is $\mathrm{G}=\operatorname{GL}^{+}(n)$, the group of invertible matrices with positive determinant in $\mathbb{R}^{n\times n}$, with Lie algebra $\mathfrak{g}=\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$. We take $\mathfrak{a}=\mathfrak{so}(n)$, the Lie algebra of antisymmetric matrices. For $\mathtt{g}\in\mathfrak{g}$, we have $\mathtt{g}_{\mathfrak{a}}=\mathtt{g}_{\mathrm{skew}}$. We typically take $\beta_{0}=1$, for $\mathtt{g}\in\mathfrak{g}$, the metric becomes
(28) $\operatorname{Tr}\mathtt{g}^{2}-(\beta+1)\operatorname{Tr}\mathtt{g}_{\mathrm{skew}}\mathtt{g}_{\mathrm{skew}}=\operatorname{Tr}\mathtt{g}^{2}_{\mathrm{sym}}+\beta\operatorname{Tr}\mathtt{g}_{\mathrm{skew}}^{\mathsf{T}}\mathtt{g}_{\mathrm{skew}}=\lvert\mathtt{g}\rvert_{\mathcal{F}}^{2}+(\beta-1)\lvert\mathtt{g}_{\mathrm{skew}}\rvert_{\mathcal{F}}^{2}$
after the orthogonal decomposition $\mathtt{g}=\mathtt{g}_{\mathrm{sym}}+\mathtt{g}_{\mathrm{skew}}$. When $\beta=1$, this is the Frobenius form, covered in [28]. Expanding the $\mathrm{skew}$ operator, theorem 1 implies
Proposition 2.
With the metric eq. 28, the Levi-Civita connection is given by
(29) $\Gamma(\xi,\eta)=-\frac{1}{2}(\xi X^{-1}\eta+\eta X^{-1}\xi)+\frac{1+\beta}{4}X([X^{-1}\xi,(X^{-1}\eta)^{\mathsf{T}}]+[X^{-1}\eta,(X^{-1}\xi)^{\mathsf{T}}]).$
The geodesic starting at $X$ with initial velocity $\xi$ is given by
(30) $\gamma(t)=X\exp(\frac{t}{2}\left((1-\beta)\xi+(1+\beta)\xi^{\mathsf{T}}\right))\exp(t(1+\beta)\xi_{\mathrm{skew}}).$
Parallel transport of a tangent vector $\eta$ at $X$ along $\gamma(t)$ is given by
(31) $\mathtt{T}_{0}^{t}(\gamma)\eta=X\exp(\frac{t}{2}\left((1-\beta)\xi+(1+\beta)\xi^{\mathsf{T}}\right))\operatorname{expv}(t\mathrm{P},X^{-1}\eta)\exp(t(1+\beta)(\xi_{\mathrm{skew}})),$
where the operator $\mathrm{P}=\mathrm{P}_{\mathtt{a}}$ for $\mathtt{a}=X^{-1}\xi$ is given by
(32) $\mathrm{P}_{\mathtt{a}}\mathtt{b}=\frac{1}{2}([\mathtt{b},\mathtt{a}]+(1+\beta)([\mathtt{a}_{\mathrm{skew}},\mathtt{b}]-[\mathtt{b}_{\mathrm{skew}},\mathtt{a}])).$
Instead of taking $\mathfrak{a}=\mathfrak{so}(n)$, if we take $\mathfrak{a}=\mathfrak{so}(n)\oplus\mathbb{R}\operatorname{I}_{n}$, then this violates condition item 3. However, if instead of taking a scalar $\beta_{1}$, we take the operator $\mathrm{B}_{1}$ which acts by multiplying $\mathfrak{so}(n)$ by $\beta_{1}$ and $\mathbb{R}\operatorname{I}_{n}$ by a negative constant, so that the metric in section 3.1 is Riemannian, the formulas still hold, as $\mathbb{R}\operatorname{I}_{n}$ does not contribute to the Lie brackets. This is the metric in [18].
The case $\beta=-1$ corresponds to the trace form and the pseudo-Riemannian Khvedelidze-Mladenov metric [14].
5. The Special Orthogonal group with a deformed metric
In this case, $\mathrm{G}=\mathrm{SO}(n)$, the group of orthogonal matrices with determinant $1$. Its Lie algbra $\mathfrak{g}=\mathfrak{so}(n)$ is identified with the set of antisymmetric matrices. For $\mathfrak{a}$, we take the algebra of matrices where the top $d\times d$ block is antisymmetric and the remaining entries vanish, ($\mathfrak{a}=\begin{bmatrix}\mathfrak{so}(d)&0_{d\times(n-d)}\\ 0_{(n-d)\times d}&0_{(n-d)\times(n-d)}\end{bmatrix}$ graphically). Here, the projection $\mathrm{p}_{\mathfrak{a}}$ from $\mathbb{R}^{n\times n}$ maps a matrix $\omega$ to the matrix $\omega_{\mathfrak{a}}$ with the top $d\times d$ block is antisymmetrized from the corresponding block of $\omega$, while the remaining entries vanish. The standard normalization takes $\beta_{0}=-\frac{1}{2}$. Setting $\beta_{1}=\alpha$, we have
Proposition 3.
At $X\in\mathrm{SO}(n)$, $\xi,\eta\in T_{X}\mathrm{SO}(n)$ where $\xi=X\begin{bmatrix}A_{\xi}&-B^{\mathsf{T}}_{\xi}\\ B_{\xi}&C_{\xi}\end{bmatrix}$, $\eta=X\begin{bmatrix}A_{\eta}&-B^{\mathsf{T}}_{\eta}\\ B_{\eta}&C_{\xi}\end{bmatrix}$, with $A_{\xi},A_{\eta}\in\mathfrak{so}(d),C_{\xi},C_{\eta}\in\mathfrak{so}(n-d)$, the metric in theorem 1 with $\beta_{0}=-\frac{1}{2},\beta_{1}=\alpha$ is given by
(33) $\langle\xi,\eta\rangle=\frac{1}{2}\operatorname{Tr}\xi^{\mathsf{T}}\eta+(\alpha-\frac{1}{2})\operatorname{Tr}((\xi^{\mathsf{T}}X)_{\mathfrak{a}}(X^{\mathsf{T}}\eta)_{\mathfrak{a}})=\frac{1}{2}\operatorname{Tr}\xi^{\mathsf{T}}\eta+(\alpha-\frac{1}{2})\operatorname{Tr}A_{\xi}^{\mathsf{T}}A_{\eta}.$
The Christoffel function of the Levi-Civita connection is
(34) $\begin{gathered}\Gamma(\xi,\eta)=\frac{1}{2}X(\xi^{\mathsf{T}}\eta+\eta^{\mathsf{T}}\xi)+\frac{1-2\alpha}{2}X([(X^{\mathsf{T}}\xi)_{\mathfrak{a}},X^{\mathsf{T}}\eta]+[(X^{\mathsf{T}}\eta)_{\mathfrak{a}},X^{\mathsf{T}}\xi])\\ =\frac{1}{2}X(\xi^{\mathsf{T}}\eta+\eta^{\mathsf{T}}\xi)+\frac{1-2\alpha}{2}X\begin{bmatrix}0_{d\times d}&-A_{\xi}B_{\eta}^{\mathsf{T}}-A_{\eta}B_{\xi}^{\mathsf{T}}\\ -B_{\eta}A_{\xi}-B_{\xi}A_{\eta}&0_{(n-d)\times(n-d)}.\end{bmatrix}.\end{gathered}$
The geodesic $\gamma(t)$ with $\gamma(0)=X,\dot{\gamma}(0)=\xi$ is
(35) $\gamma(t)=X\exp\left(t\begin{bmatrix}2\alpha A_{\xi}&-B_{\xi}^{\mathsf{T}}\\ B_{\xi}&C_{\xi}\end{bmatrix}\right)\exp\left(t\begin{bmatrix}(1-2\alpha)A_{\xi}&0\\ 0&0\end{bmatrix}\right).$
Parallel transport of $\eta$ along $\gamma(t)$ is given by
(36) $\mathtt{T}(\gamma)_{X}^{\mathsf{T}}\eta=X\exp\left(t\begin{bmatrix}2\alpha A_{\xi}&-B^{\mathsf{T}}_{\xi}\\ B_{\xi}&C_{\xi}\end{bmatrix}\right)\operatorname{expv}(t\mathrm{P},X^{\mathsf{T}}\eta)\exp\left(t\begin{bmatrix}(1-2\alpha)A_{\xi}&0\\ 0&0\end{bmatrix}\right)$
where the operator $\mathrm{P}=\mathrm{P}_{\mathtt{a}}$ for $\mathtt{a}=X^{\mathsf{T}}\xi$ is given in eq. 21 with $\beta=-2\alpha$.
6. A quotient metric
The two standard metrics on the Stiefel manifold are instances of the quotient metric defined in this section. This metric takes advantage of the fact that the trace form is bi-invariant, to construct a right-invariant group action of a simple format.
Let $\mathrm{K}\subset\mathrm{G}$ be a connected Lie subgroup. We have an action of $\mathrm{K}$ on $\mathrm{G}$ by right multiplication, the action is $(X,K)\mapsto XK\in\mathrm{G}$, for $X\in\mathrm{G},K\in\mathrm{K}$. We assume the action is proper (a condition that is satisfied if $\mathrm{K}$ is compact, that is the case for the examples we consider here). Then $\mathcal{M}=\mathrm{G}/\mathrm{K}$ has the structure of a quotient manifold. We will show in theorem 2 below that with an additional condition, $\mathrm{K}$ acts on $\mathrm{G}$ as an isometry group of the metric in theorem 1 by right multiplication, and we can apply the theory of pseudo-Riemannian submersion.
Theorem 2.
Assume the connected subgroup $\mathrm{K}=\mathrm{K}_{\mathfrak{a}}\times\mathrm{K}_{\mathfrak{a}_{{}_{\top}}}\subset\mathrm{G}$ with Lie algera $\mathfrak{k}=\mathfrak{k}_{\mathfrak{a}}\oplus\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}$ satisfies
(37) $\mathfrak{k}_{\mathfrak{a}}\subset\mathfrak{a};\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}\subset\mathfrak{a}_{{}_{\top}}\text{ and }\mathfrak{k}^{\mathsf{T}}=\mathfrak{k}$
and $\mathrm{K}$ acts smoothly and properly on $\mathrm{G}$ by right multiplication. Then $\mathrm{K}$ acts as isometries of the metric in theorem 1, and $\mathcal{M}=\mathrm{G}/\mathrm{K}$ can be given a quotient metric. Let $\mathfrak{m}=\mathfrak{k}_{{}_{\perp}}$. For $X\in\mathrm{G}$, the vertical space $\mathcal{V}_{X}$ could be identified with $X\mathfrak{k}$, the horizontal space could be identified with $X\mathfrak{m}$. For two horizontal vector fields $\mathtt{Y},\mathtt{Z}$, the horizontal connection is given by $\nabla^{\mathcal{H}}_{\mathtt{Y}}\mathtt{Z}=\mathrm{D}_{\mathtt{X}}\mathtt{Y}+\Gamma^{\mathcal{H}}(\mathtt{X},\mathtt{Y})$, where the horizontal Christoffel function $\Gamma^{\mathcal{H}}$ for two horizontal vectors $\xi,\eta$ is given by
(38) $\begin{gathered}\Gamma^{\mathcal{H}}(\xi,\eta)=-\frac{1}{2}\left(\xi X^{-1}\eta+\eta X^{-1}\xi+X[X^{-1}\xi,X^{-1}\eta]_{\mathfrak{k}}\right)\\ +\frac{1+\beta}{2}X\left([(X^{-1}\xi)_{\mathfrak{a}},X^{-1}\eta]+[(X^{-1}\eta)_{\mathfrak{a}},X^{-1}\xi]\right).\end{gathered}$
The horizontal geodesic $\gamma$ with $\gamma(0)=X,\dot{\gamma}(0)=\xi$, where $\xi$ is a horizontal vector is given by eq. 20. Let $\mathtt{a}=X^{-1}\xi$. The parallel transport of a horizontal vector $\eta\in T_{X}\mathcal{M}$ is given by
(39) $\begin{gathered}\mathtt{T}(\gamma)_{0}^{t}\eta=X\operatorname{expm}(t\left(-(1+\beta)\mathtt{a}_{\mathfrak{a}}+\mathtt{a}\right))W(t)U(t)\end{gathered}$
where $U(t):=\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}})$ and $W(t)$ is a solution of the equation $W(0)=X^{-1}\eta$,
(40) $\dot{W}=\frac{1}{2}\left([W,\mathtt{a}]-U(U^{-1}[W,\mathtt{a}]U)_{\mathfrak{k}}U^{-1}+(1+\beta)([\mathtt{a}_{\mathfrak{a}},W]-[W_{\mathfrak{a}},\mathtt{a}])\right).$
Assume further that $(U^{-1}[W,\mathtt{a}]U)_{\mathfrak{k}}=U^{-1}[W,\mathtt{a}]_{\mathfrak{k}}U$, then $W=\operatorname{expv}(\mathrm{P}_{\mathtt{a}},X^{-1}\eta)$ where for $\mathtt{b}\in\mathfrak{g}$
(41) $\mathrm{P}_{\mathtt{a}}:\mathtt{b}\mapsto\frac{1}{2}\left([\mathtt{b},\mathtt{a}]_{\mathfrak{m}}+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}])\right).$
This condition is satisfied if $\beta=-1$ or $\mathfrak{k}\subset\mathfrak{a}_{{}_{\top}}$. In these two cases, $\mathrm{P}_{\mathfrak{a}}$ maps $\mathfrak{m}$ to itself and the operator $\mathrm{P}_{\mathfrak{a}}$ is antisymmetric in the pairing $\langle\rangle_{\vec{\beta}}$ on $\mathfrak{m}$.
We note the horizontal Christoffel function differs from the tangent one eq. 18 by a vertical term $\frac{1}{2}X[X^{-1}\xi,X^{-1}\eta]_{\mathfrak{k}}$.
We will see the condition for the simplified parallel transport formula is satisfied for Stiefel manifolds.
Proof.
As $\mathfrak{a}$ and $\mathfrak{a}_{{}_{\top}}$ commute, $\mathrm{K}_{\mathfrak{a}}$ and $\mathrm{K}_{\mathfrak{a}_{{}_{\top}}}$ also commute, thus, we only need to prove each subgroup acts as isometries. For $\mathtt{k}\in\mathfrak{k}$, set $K=\exp(\mathtt{k})\in\mathrm{K}$. For $\eta=X\mathtt{g}$ in $T_{X}\mathrm{G}$ with $\mathtt{g}\in\mathfrak{g}$, right multiplication by $K$ sends $X$ to $XK$ and $\eta$ to $\eta K$. Hence,
$\begin{gathered}\langle\eta K,\eta K\rangle_{XK,\vec{\beta}}=\langle(XK)^{-1}X\mathtt{g}K,(XK)^{-1}X\mathtt{g}K\rangle_{\vec{\beta}}\\ =\beta_{0}\operatorname{Tr}(K^{-1}\mathtt{g}K)^{2}-\beta_{0}(1+\beta)\operatorname{Tr}((K^{-1}\mathtt{g}K)_{\mathfrak{a}})^{2}.\end{gathered}$
Note $K^{-1}\mathtt{g}K=\operatorname{expv}(-\operatorname{ad}_{\mathtt{k}},\mathtt{g})$. If $\mathtt{k}\in\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}$ then $\operatorname{ad}_{\mathtt{k}}\mathfrak{g}\subset\mathfrak{a}_{{}_{\perp}}$. Hence, $(\operatorname{ad}_{\mathtt{k}}^{j}\mathtt{g})_{\mathfrak{a}}=0$ for $j>0$, and $(K^{-1}\mathtt{g}K)_{\mathfrak{a}}=\mathtt{g}_{\mathfrak{a}}$. Thus, $\mathrm{K}_{\mathfrak{a}_{{}_{\top}}}$ acts as isometries, as the above reduces to
$\beta_{0}\operatorname{Tr}K^{-1}\mathtt{g}^{2}K-\beta_{0}(1+\beta)\operatorname{Tr}\mathtt{g}_{\mathfrak{a}}^{2}=\beta_{0}\operatorname{Tr}\mathtt{g}^{2}-\beta_{0}(1+\beta)\operatorname{Tr}\mathtt{g}_{\mathfrak{a}}^{2}.$
Similarly, if $\mathtt{k}\in\mathfrak{k}_{\mathfrak{a}}$ then $\operatorname{ad}_{\mathtt{k}}^{j}\mathtt{g}\subset\operatorname{ad}_{\mathtt{k}}^{j}\mathtt{g}_{\mathfrak{a}}+\mathfrak{a}_{{}_{\Join}}$ for $j>0$ by induction, and $(K^{-1}\mathtt{g}K)_{\mathfrak{a}}=K^{-1}\mathtt{g}_{\mathfrak{a}}K$. Then $\langle\eta K,\eta K\rangle_{XK,\vec{\beta}}$ reduces again to $\beta_{0}\operatorname{Tr}\mathtt{g}^{2}-\beta_{0}(1+\beta)\operatorname{Tr}\mathtt{g}_{\mathfrak{a}}^{2}$. Hence, $\mathrm{K}_{\mathfrak{a}}$ and thus $\mathrm{K}$ are groups of isometries.
We need to prove for two horizontal vector fields $\mathtt{Y},\mathtt{Z}$, $\nabla^{\mathcal{H}}_{\mathtt{Y}}\mathtt{Z}$ with $\Gamma^{\mathcal{H}}$ given in eq. 38 is horizontal, metric compatible and projects to a torsion-free connection in the quotient. The latter is true as the only nonsymmetric term $\frac{1}{2}X[X^{-1}\xi,X^{-1}\eta]_{\mathfrak{k}}$ is vertical. The second line in eq. 38 is in $\mathfrak{a}_{{}_{\Join}}$, thus, is horizontal since $\mathfrak{k}\subset\mathfrak{a}\oplus\mathfrak{a}_{{}_{\top}}$. If $\mathtt{Y}(X)$ and $\mathtt{Z}$ are horizontal, with $\mathtt{Y}(X)=X\mathtt{y}(X),\mathtt{Z}(X)=X\mathtt{z}(x)$ for $\mathtt{y}(X),\mathtt{z}(X)\in\mathfrak{m}$ then
$\begin{gathered}\mathrm{D}_{\mathtt{Y}}\mathtt{Z}-\frac{1}{2}\left((\mathtt{Y}X^{-1}\mathtt{Z}+\mathtt{Z}X^{-1}\mathtt{Y})+X[X^{-1}\mathtt{Y},X^{-1}\mathtt{Z}]_{\mathfrak{k}}\right)\\ =X\mathtt{y}\mathtt{z}+X\mathrm{D}_{\mathtt{Y}}\mathtt{z}-\frac{1}{2}\left((X\mathtt{y}\mathtt{z}+X\mathtt{z}\mathtt{y})+X[\mathtt{y},\mathtt{z}]_{\mathfrak{k}}\right)\\ =X\mathrm{D}_{\mathtt{Y}}\mathtt{z}+\frac{1}{2}X[\mathtt{y},\mathtt{z}]-\frac{1}{2}X[\mathtt{y},\mathtt{z}]_{\mathfrak{k}}=X\mathrm{D}_{\mathtt{Y}}\mathtt{z}+\frac{1}{2}X[\mathtt{y},\mathtt{z}]_{\mathfrak{m}}.\end{gathered}$
The last expression is horizontal because $\mathrm{D}_{\mathtt{Y}}\mathtt{z}$ is horizontal by linearity. Finally, metric compatibility with horizontal vector fields follows since the extra term $\frac{1}{2}X[\mathtt{y},\mathtt{z}]_{\mathfrak{k}}$ is vertical, which does not contribute when paired with a horizontal vector.
That the horizontal geodesic is a geodesic with initial horizontal velocity is well-known, or we can check the formula directly, as $\frac{1}{2}\gamma[\gamma^{-1}\dot{\gamma},\gamma^{-1}\dot{\gamma}]_{\mathfrak{k}}$ vanishes.
Let $\Delta(t)$ be the transport of $\eta$ along $\gamma(t)$, set $\Delta=\gamma U^{-1}WU$, or $W(t)=U(t)\gamma(t)^{-1}\Delta(t)U(t)^{-1}$, we derive its equation as in theorem 1
$\begin{gathered}\dot{\Delta}=\gamma U^{-1}(\mathtt{a}W+(1+\beta)[W,\mathtt{a}_{\mathfrak{a}}]+\dot{W})U,\\ \Gamma(\dot{\gamma},\Delta)=-\frac{1}{2}\gamma U^{-1}(\mathtt{a}W+W\mathtt{a}+U[U^{-1}\mathtt{a}U,U^{-1}WU]_{\mathfrak{k}}U^{-1})U\\ +\frac{1+\beta}{2}\gamma U^{-1}\left([\mathtt{a}_{\mathfrak{a}},W]+[W_{\mathfrak{a}},\mathtt{a}]\right)U,\end{gathered}$
which simplifies to the equation for $W$. When $(U^{-1}[W,\mathtt{a}]U)_{\mathfrak{k}}=U^{-1}[W,\mathtt{a}]_{\mathfrak{k}}U$ holds, the coefficients of that equation are constant and we have eq. 41.
That simplified condition holds when $\beta=-1$, as it implies $U=\operatorname{I}_{n}$ and the rest follows. In this case, $\mathrm{P}_{\mathtt{a}}\mathtt{b}=\frac{1}{2}[\mathtt{b},\mathtt{a}]_{\mathfrak{m}}\in\mathfrak{m}$, or $\mathrm{P}_{\mathtt{a}}$ maps to $\mathfrak{m}$.
For the remaining simplified case, when $\mathfrak{k}\subset\mathfrak{a}_{{}_{\top}}$ then we have $[\mathtt{a}_{\mathfrak{a}},\mathtt{c}]_{\mathfrak{k}}=0$ for $\mathtt{c}\in\mathfrak{g}$, since set $\mathtt{c}=\mathtt{c}_{\mathfrak{a}_{{}_{\top}}}+\mathtt{c}_{\mathfrak{a}}+\mathtt{c}_{\mathfrak{a}_{{}_{\Join}}}$, the bracket reduces to $[\mathtt{a}_{\mathfrak{a}},\mathtt{c}_{\mathfrak{a}}+\mathtt{c}_{\mathfrak{a}_{{}_{\Join}}}]_{\mathfrak{k}}\subset(\mathfrak{a}+\mathfrak{a}_{{}_{\Join}})_{\mathfrak{k}}=0$. This shows $(\operatorname{ad}_{\mathtt{a}_{\mathfrak{a}}}^{j}\mathtt{c})_{\mathfrak{k}}=0$ for $j>0$ or $(U^{-1}[W,\mathtt{a}]U)_{\mathfrak{k}}=[W,\mathtt{a}]_{\mathfrak{k}}$. Similarly, $U[W,\mathtt{a}]_{\mathfrak{k}}U^{-1}=[W,\mathtt{a}]_{\mathfrak{k}}$ as $\mathtt{a}_{\mathfrak{a}}$ commutes with $\mathfrak{k}\subset\mathfrak{a}_{{}_{\top}}$, hence the simplified condition is satisfied.
To show $\mathrm{P}_{\mathtt{a}}$ operates on $\mathfrak{m}$, the previous paragraph showed $[\mathtt{a}_{\mathfrak{a}},\mathtt{b}]_{\mathfrak{k}}=0$ and $[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]_{\mathfrak{k}}=0$, so all three terms of $\mathrm{P}_{\mathtt{a}}$ are in $\mathfrak{m}$.
By theorem 1, $\mathrm{P}^{\mathfrak{g}}_{\mathtt{a}}:\mathtt{b}\mapsto\frac{1}{2}\left([\mathtt{b},\mathtt{a}]+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}])\right)$ is antisymmetric under the inner product $\langle\rangle_{\vec{\beta}}$. Thus, we only need to prove $\mathtt{b}\mapsto[\mathtt{b},\mathtt{a}]_{\mathfrak{k}}$ is antisymmetric, when restricted to $\mathfrak{m}$. But this is clear as $\mathfrak{k}$ and $\mathfrak{m}$ are orthogonal under $\langle\rangle_{\vec{\beta}}$. ∎
7. Stiefel manifolds
The Stiefel manifold $\mathrm{St}_{n,d}$ consists of orthogonal matrices in $\mathbb{R}^{n\times d}$, where we assume $d<n$. Thus, $Y\in\mathrm{St}_{n,d}$ if $Y^{\mathsf{T}}Y=\operatorname{I}_{d}$. The simplest Stiefel manifolds are unit spheres, corresponding to the case $d=1$. The tangent space to $\mathrm{St}_{n,d}$ at $Y$ is defined by the equation $(Y^{\mathsf{T}}\xi)_{\mathrm{sym}}=0$, $\xi\in\mathbb{R}^{n\times d}$.
7.1. An isometry and a totally geodesic submanifold
In a sphere, great circles are geodesics. Given a point $Y$ on the unit sphere, and a tangent vector $\xi$, the plane spanned by $Y$ and $\xi$ cuts a great circle. Changing to the coordinate defined by $Y$ and the unit vector in direction $\xi$, this great circle maps bijectively and isometrically to the unit circle in $\mathbb{R}^{2}$. There is a similar picture for Stiefel manifolds. We start the following lemma, first discussed in [19, section 5.1].
Lemma 2.
Let $Y$ be a point on the Stiefel manifold $\mathcal{M}=\mathrm{St}_{n,d}\subset\mathbb{R}^{n\times d}$, and $\xi$ be a tangent vector to $\mathcal{M}$ at $Y$. If $\xi\neq YY^{\mathsf{T}}\xi$ and $Q_{\xi}$ is an orthonormal basis of the column span of $\xi-YY^{\mathsf{T}}\xi$, then $\xi=YA+Q_{\xi}R_{\xi}$ for some matrix $R_{\xi}$. More generally, assume $Q$ is an orthogonal matrix in $\mathbb{R}^{n\times k}$ for an integer $k\geq 0$, such that $Y^{\mathsf{T}}Q=0$ and $\xi=YA+QR$ for a matrix $R\in\mathbb{R}^{k\times d}$. Then the vector subspace $\mathcal{S}=\mathcal{S}_{Q}\subset\mathbb{R}^{n\times d}$ of matrices of the form
(42) $\mathcal{S}:=\{V\in\mathbb{R}^{n\times d}|\;V=YV_{Y}+QV_{Q}\text{ for }V_{Y}\in\mathbb{R}^{d\times d},V_{Q}\in\mathbb{R}^{k\times d}\}$
intersects $\mathrm{St}_{n,d}$ along a submanifold $\mathcal{M}_{\mathcal{S}}\subset\mathcal{M}$, which maps bijectively to $\mathrm{St}_{d+k,d}$. This bijection could be defined as
(43) $\mathfrak{F}:V\mapsto F^{\mathsf{T}}V\text{ for }F=F_{Y,Q}=\begin{bmatrix}Y|Q\end{bmatrix},\text{ with }V\in\mathcal{M}_{\mathcal{S}}.$
We have $\mathfrak{F}Y=\operatorname{I}_{d+k,d}=\begin{bmatrix}\operatorname{I}_{d}\\ O_{k\times d}\end{bmatrix}\in\mathrm{St}_{d+k,d}\subset\mathbb{R}^{n\times(d+k)}$. If $\xi=YY^{\mathsf{T}}\xi$, we can consider $F_{Y,Q}=Y$, or $Q=0$ and the intersection $\mathcal{M}_{\mathcal{S}}$ is isometric to $\operatorname{O}(d)$, with two connected components.
Proof.
First, if $\xi\neq YY^{\mathsf{T}}\xi$, since columns of $Q_{\xi}$ are in the column span of $\xi-YY^{\mathsf{T}}\xi$, we have $Y^{\mathsf{T}}Q_{\xi}=0$, and $Q_{\xi}^{\mathsf{T}}Q_{\xi}=\operatorname{I}_{k}$ by the orthogonal assumption. Express the columns of $\xi-YY^{\mathsf{T}}\xi$ in the basis $Q_{\xi}$ as $\xi-YY^{\mathsf{T}}\xi=Q_{\xi}R_{\xi}$ for some matrix $R_{\xi}$, then $\xi=Y(Y^{\mathsf{T}}\xi)+Q_{\xi}R_{\xi}$. This shows $Q_{\xi}$ has the desired property.
Assume $Y^{\mathsf{T}}Q=0,Q^{\mathsf{T}}Q=\operatorname{I}_{k}$ and $\xi=YA+QR$ for some matrix $A,R$. With $F=\begin{bmatrix}Y|Q\end{bmatrix}$, we have $F^{\mathsf{T}}F=\operatorname{I}_{d+k,d+k}$. If $V=YV_{Y}+QV_{Q}\in\mathcal{M}_{\mathcal{S}}$, then $\mathfrak{F}V=F^{\mathsf{T}}V=\begin{bmatrix}V_{Y}\\ V_{Q}\end{bmatrix}$. The condition $V^{\mathsf{T}}V=\operatorname{I}_{d}$ implies $V_{Y}^{\mathsf{T}}V_{Y}+V_{Q}^{\mathsf{T}}V_{Q}=\operatorname{I}_{d}$ by substitution, but this means $(\mathfrak{F}V)^{\mathsf{T}}(\mathfrak{F}V)=\operatorname{I}_{d}$. Thus, $\mathfrak{F}$ maps to $\mathrm{St}_{d+k,d}$. Furthermore, $V=F(\mathfrak{F}V)$ allows us to compute the inverse of $\mathfrak{F}$ to $\mathcal{M}_{\mathcal{S}}$, hence $\mathfrak{F}$ is a bijection.
By direct calculation, $F^{\mathsf{T}}Y=\operatorname{I}_{d+k,d}$ and $F^{\mathsf{T}}\xi=\begin{bmatrix}A\\ R\end{bmatrix}\in T_{\operatorname{I}_{d+k,d}}\mathrm{St}_{d+k,d}$.
The case $k=0$ follows similarly. ∎
We will see the metric induced from the family in section 5 could be expressed as
(44) $\langle\xi,\xi\rangle_{\vec{\beta}}=\operatorname{Tr}\xi^{\mathsf{T}}\xi+(\alpha-1)\operatorname{Tr}\xi^{\mathsf{T}}YY^{\mathsf{T}}\xi,$
which is the metric [11], using the parametrization in [19]. For a tangent vector $\nu$ at $V\in\mathcal{M}_{\mathcal{S}}$, write $\nu=F\mathfrak{F}(\nu),V=F\mathfrak{F}(V)$ and substitute, we get
$\langle\nu,\nu\rangle_{\vec{\beta}}=\operatorname{Tr}\mathfrak{F}(\nu)^{\mathsf{T}}\mathfrak{F}(\nu)+(\alpha-1)\operatorname{Tr}\mathfrak{F}(\nu)^{\mathsf{T}}\mathfrak{F}(V)\mathfrak{F}(V)^{\mathsf{T}}\mathfrak{F}(\nu).$
Thus, if we equip $\mathrm{St}_{d+k,d}$ with a metric with the same parameter $\alpha$, $\mathfrak{F}$ is an isometry.
Computation done on this smaller Stiefel manifold is more effective, as observed in the case of geodesics [5, 19, 31]. The case of parallel transport will be done similarly.
By [23, page 104], $\mathcal{M}_{\mathcal{S}}$ is a totally geodesic submanifold, as the geodesic with initial point $V\in\mathcal{M}_{\mathcal{S}}$ and initial vector tangent to $\mathcal{M}_{\mathcal{S}}$ will stay on $\mathcal{M}_{\mathcal{S}}$.
7.2. Submersion metric, connection and parallel transport
The Stiefel manifold $\mathrm{St}_{n,d}$ could be consider as a quotient manifold of $\mathrm{SO}(n)$ under the map sending an $n\times n$ matrix to its first $d$ columns, or $\mathfrak{q}:X\mapsto X\operatorname{I}_{n,d}$, for $X\in\mathrm{SO}(n)$, $\operatorname{I}_{n,d}=\begin{bmatrix}\operatorname{I}_{d}\\ 0_{(n-d)\times d}\end{bmatrix}$. This map is onto, realizing $\mathrm{St}_{n,d}$ as a quotient of $\mathrm{SO}(n)$ by the group $\mathrm{K}$ of block diagonal matrices of the form $\operatorname{diag}(\operatorname{I}_{d},Q)$ where $Q\in\mathrm{SO}(n-d)$. This is due to the fact that we can complement an orthogonal matrix $Y$ with an orthogonal matrix $Y_{\perp}\in\mathrm{St}_{n,n-d}$ to an element of $\mathrm{SO}(n)$, and given $Y_{\perp}$, any other complement basis is of the form $Y_{\perp}Q$ for $Q\in\mathrm{K}$. In this case, $\mathfrak{a}$ is the top $d\times d$ block-matrix copy of $\mathfrak{so}(d)$ and $\mathfrak{k}=\mathfrak{a}_{{}_{\top}}$ is the bottom $(n-d)\times(n-d)$ copy of $\mathfrak{so}(n-d)$. The complement $\mathfrak{m}$ consists of all matrices with the lower right $(n-d)\times(n-d)$ blocks identically zero. Thus, graphically, $\mathfrak{k}=\begin{bmatrix}0_{d\times d}&0\\ 0&\mathfrak{so}(n-d)\end{bmatrix},\mathfrak{m}=\begin{bmatrix}\mathfrak{a}&\mathfrak{a}_{{}_{\Join}}\\ \mathfrak{a}_{{}_{\Join}}&0_{(n-d)\times(n-d)}\end{bmatrix}$.
The differential $d\mathfrak{q}$ at $\operatorname{I}_{n}$ also maps a Lie algebra element in $\mathfrak{so}(n)$ to its first $d$ columns. If $\Delta=\begin{bmatrix}A\\ B\end{bmatrix}$ is a tangent vector at $\operatorname{I}_{n,d}$, then its horizontal lift is the matrix $\bar{\Delta}=\begin{bmatrix}A&-B^{\mathsf{T}}\\ B&0\end{bmatrix}$, the bottom right entry corresponding to $\mathfrak{k}$ is set to zero. With $\beta_{0}=-\frac{1}{2},\beta_{1}=\alpha,\beta=\frac{\beta_{1}}{\beta_{0}}=-2\alpha$, the metric in eq. 33 at $\operatorname{I}_{n}$ is
$\langle\bar{\Delta},\bar{\Delta}\rangle_{\vec{\beta}}=\frac{1}{2}(\operatorname{Tr}A^{\mathsf{T}}A+2\operatorname{Tr}B^{\mathsf{T}}B)+(\alpha-\frac{1}{2})\operatorname{Tr}A^{\mathsf{T}}A=\alpha\operatorname{Tr}A^{\mathsf{T}}A+\operatorname{Tr}B^{\mathsf{T}}B.$
More generally, for $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\in\mathrm{SO}(n)$, split $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu=\begin{bmatrix}Y|Y_{{}_{\perp}}\end{bmatrix}$ along the $d$-th column, if $\bar{\xi}\in T_{\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu}\mathrm{SO}(n)$ is horizontal and maps to $\xi\in T_{Y}\mathrm{St}_{n,d}$, with $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu^{\mathsf{T}}\bar{\xi}=\bar{\Delta}$ above then we deduce $A=Y^{\mathsf{T}}\xi,B=Y_{{}_{\perp}}^{\mathsf{T}}\xi$. Using $Y_{{}_{\perp}}Y_{{}_{\perp}}^{\mathsf{T}}+YY^{\mathsf{T}}=\operatorname{I}_{n}$, the metric is as given by eq. 44
$\alpha\operatorname{Tr}\xi YY^{\mathsf{T}}\xi+\operatorname{Tr}\xi Y_{{}_{\perp}}Y_{{}_{\perp}}^{\mathsf{T}}\xi=\operatorname{Tr}\xi^{\mathsf{T}}\xi+(\alpha-1)\operatorname{Tr}\xi^{\mathsf{T}}YY^{\mathsf{T}}\xi.$
Moreover, set $\operatorname{I}_{{}_{\perp}}=\begin{bmatrix}0_{d\times(n-d)}\\ \operatorname{I}_{n-d}\end{bmatrix}$ then the horizontal lift $\bar{\xi}$ is given by
(45) $\bar{\xi}=\xi\operatorname{I}_{n,d}^{\mathsf{T}}-Y\xi^{\mathsf{T}}Y_{{}_{\perp}}\operatorname{I}_{{}_{\perp}}^{\mathsf{T}}.$
To see this, with $\bar{\xi}$ as above, we have $\bar{\xi}\operatorname{I}_{n,d}=\xi$ and using $Y^{\mathsf{T}}Y=\operatorname{I}_{d},Y_{{}_{\perp}}^{\mathsf{T}}Y=0$ then
$\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu^{\mathsf{T}}\bar{\xi}=(Y\operatorname{I}_{n,d}^{\mathsf{T}}+Y_{{}_{\perp}}\operatorname{I}_{{}_{\perp}}^{\mathsf{T}})^{\mathsf{T}}\bar{\xi}=\operatorname{I}_{n,d}Y^{\mathsf{T}}\xi\operatorname{I}_{n,d}-\operatorname{I}_{n,d}\xi^{\mathsf{T}}Y_{{}_{\perp}}\operatorname{I}_{{}_{\perp}}^{\mathsf{T}}+\operatorname{I}_{{}_{\perp}}Y_{{}_{\perp}}^{\mathsf{T}}\xi\operatorname{I}_{n,d}^{\mathsf{T}},$
which is antisymmetric, and $\operatorname{I}_{{}_{\perp}}^{\mathsf{T}}(X^{\mathsf{T}}\bar{\xi})\operatorname{I}_{{}_{\perp}}=0$. We also have $(X^{\mathsf{T}}\bar{\xi})_{\mathfrak{a}}=\operatorname{I}_{n,d}Y^{\mathsf{T}}\xi\operatorname{I}_{n,d}$.
Under this metric and with the lift above, using eq. 38, we can show [11, 20] the Levi-Civita connection for Stiefel manifold is $\nabla_{\mathtt{Z}}\mathtt{W}=\mathrm{D}_{\mathtt{Z}}\mathtt{W}+\Gamma(\mathtt{Z},\mathtt{W})$ with the Christofel function evaluated at $Y\in\mathcal{M}=\mathrm{St}_{n,d},\mathtt{Z}(Y)=\xi,\mathtt{W}(Y)=\eta$ is
(46) $\Gamma(\xi,\eta)=\frac{1}{2}Y(\xi^{\mathsf{T}}\eta+\eta^{\mathsf{T}}\xi)+(1-\alpha)(\operatorname{I}_{n}-YY^{\mathsf{T}})(\xi\eta^{\mathsf{T}}+\eta\xi^{\mathsf{T}})Y.$
For $Y\in\mathcal{M}=\mathrm{St}_{n,d}$, with $\mathfrak{q}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)=Y$, and $\xi\in T_{Y}\mathcal{M}=T_{Y}\mathrm{St}_{n,p}$, using eq. 20, the geodesics, with initial conditions $\gamma(0)=Y,\dot{\gamma}(0)=\xi$ is
(47) $\gamma(t)=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\operatorname{expm}(t\begin{bmatrix}2\alpha A&-B^{\mathsf{T}}\\ B&0\end{bmatrix})\operatorname{expm}(t\begin{bmatrix}(1-2\alpha)A&0\\ 0&0\end{bmatrix})\operatorname{I}_{n,d},$
if $\bar{\xi}=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\begin{bmatrix}A&-B^{\mathsf{T}}\\ B&0\end{bmatrix}$. Parallel transport could also be computed using eq. 39, noting in this case, $(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu^{\mathsf{T}}\xi)_{\mathfrak{a}}=\begin{bmatrix}A&0\\ 0&0\end{bmatrix}$ commutes with $\mathfrak{k}$, and the simplified condition is satisfied. The operations are $O(n^{3})$ and not efficient. The formula is dependent on $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$.
However, we can use the isometry in section 7.1. For geodesics, the isometry gives us the geodesic formula $\gamma(t)=F\gamma_{\mathfrak{F}\mathcal{M}_{\mathcal{S}}}(t)$, where $\gamma_{\mathfrak{F}\mathcal{M}_{\mathcal{S}}}$ is the geodesic on $\mathrm{St}_{d+k,d}$ starting at $\operatorname{I}_{d+k,d}$ with initial velocity $\begin{bmatrix}A\\ R\end{bmatrix}$. We lift $\gamma_{\mathfrak{F}\mathcal{M}_{\mathcal{S}}}$ to $\mathrm{SO}(d+k)$ starting at $\operatorname{I}_{d+k}$, with initial horizontal velocity $\begin{bmatrix}A&-R^{\mathsf{T}}\\ R&0\end{bmatrix}$. The result is the known geodesic formula in [5, 31, 19]
(48) $\gamma(t)=\begin{bmatrix}Y|Q\end{bmatrix}\operatorname{expm}(t\begin{bmatrix}2\alpha A&-R^{\mathsf{T}}\\ R&0\end{bmatrix})\begin{bmatrix}\operatorname{expm}((1-2\alpha)tA)\\ 0_{k,d}\end{bmatrix}.$
The parallel transport equation eq. 39 could be used for tangent vectors of $\mathcal{M}_{\mathcal{S}}$. Since parallel transports are linear in initial vectors, for the complete picture, we will need a formula for tangent vectors to $\mathrm{St}_{n,d}$ normal to $T_{Y}\mathcal{M}_{\mathcal{S}}$. We have
Lemma 3.
For a geodesic $\gamma(t)$ with $\gamma(0)=Y\in\mathcal{M}=\mathrm{St}_{n,p},\dot{\gamma}(0)=\xi\in T_{Y}\mathcal{M}$, and $\mathcal{S}$ is the vector subspace of $\mathbb{R}^{n\times d}$ of matrices with columns in the span of columns of $Y$ and $\xi$ as in eq. 42. If $\eta\in T_{Y}\mathrm{St}_{n,d}$ satisfies both $\eta^{\mathsf{T}}Y=0$ and $\eta^{\mathsf{T}}\xi=0$, or equivalently, $\eta$ is normal to $\mathcal{S}$ in the metric eq. 44, then with $A=Y^{\mathsf{T}}\xi$, the parallel transport of $\eta$ along $\gamma$ in that metric is given by
(49) $\mathtt{T}(\gamma)_{0}^{t}\eta=\Delta(t):=\eta\operatorname{expm}(t(1-\alpha)A).$
Proof.
First, we show the condition $\eta$ is normal to $\mathcal{S}$ is equivalent to $Y^{\mathsf{T}}\eta=0$ and $\xi^{\mathsf{T}}\eta=0$. We need to show $\langle\eta,YA_{1}+QR_{1}\rangle_{\vec{\beta}}=0$ for all anti symmetric $A_{1}$ and $R_{1}\in\mathbb{R}^{k\times d}$ means $\eta^{\mathsf{T}}Y=0$ and $\eta^{\mathsf{T}}Q=0$. But $\langle\eta,YA_{1}\rangle_{\vec{\beta}}=\alpha\operatorname{Tr}\eta^{\mathsf{T}}YA_{1}$ and $\langle\eta,QR_{1}\rangle_{\vec{\beta}}=\operatorname{Tr}\eta^{\mathsf{T}}QR_{1}$ and $\eta^{\mathsf{T}}Y$ is antisymmetric, hence $\eta^{\mathsf{T}}Y=0,\eta^{\mathsf{T}}Q=0$. The reverse direction is clear.
The parallel transport equation is $\dot{\Delta}+\Gamma(\gamma(t);\dot{\gamma}(t),\Delta(t))=0$. With $\gamma$ in eq. 48
$\dot{\gamma}(t)=\begin{bmatrix}Y|Q\end{bmatrix}\operatorname{expm}(t\begin{bmatrix}2\alpha A&-R^{\mathsf{T}}\\ R&0\end{bmatrix})\begin{bmatrix}A&-R^{\mathsf{T}}\\ R&0\end{bmatrix}\begin{bmatrix}\operatorname{expm}((1-2\alpha)tA)\\ 0_{k,d}\end{bmatrix}.$
and assuming eq. 49, $\dot{\Delta}=(1-\alpha)\eta e^{t(1-\alpha)A}A$. Write $\gamma=\begin{bmatrix}Y|Q\end{bmatrix}C_{1},\dot{\gamma}=\begin{bmatrix}Y|Q\end{bmatrix}C_{2}$ then $\Delta^{\mathsf{T}}\gamma$ and $\Delta^{\mathsf{T}}\dot{\gamma}$ simplifies to
$\begin{gathered}\Delta^{\mathsf{T}}\gamma=(e^{-t(1-\alpha)A})\eta^{\mathsf{T}}\begin{bmatrix}Y|Q\end{bmatrix}C_{1}=0,\\ \Delta^{\mathsf{T}}\dot{\gamma}=(e^{-t(1-\alpha)A})\eta^{\mathsf{T}}\begin{bmatrix}Y|Q\end{bmatrix}C_{2}=0\end{gathered}$
by the initial condition. Thus, $\Gamma(\dot{\gamma},\Delta)=(1-\alpha)\Delta\dot{\gamma}^{\mathsf{T}}\gamma$ and
$\begin{gathered}\dot{\Delta}+\Gamma(\dot{\gamma},\Delta)=\dot{\Delta}+(1-\alpha)\Delta\begin{bmatrix}\operatorname{expm}((1-2\alpha)tA)\\ 0_{k,d}\end{bmatrix}^{\mathsf{T}}\begin{bmatrix}-A&R^{\mathsf{T}}\\ -R&0\end{bmatrix}\begin{bmatrix}\operatorname{expm}((1-2\alpha)tA)\\ 0_{k,d}\end{bmatrix}\\ =\dot{\Delta}-(1-\alpha)\Delta A=0\end{gathered}$
which is satisfied by $\Delta(t)$ in eq. 49. ∎
Putting everything together, we have
Theorem 3.
For $Y\in\mathrm{St}_{n,d}$ and $\xi\in T_{Y}\mathrm{St}_{n,d}$, let $k\leq d$ be the rank of the column span of $\xi-YY^{\mathsf{T}}\xi$, and let $Q$ be an orthonormal basis of that span, take $Q=0$ if $\xi-YY^{\mathsf{T}}\xi=0$. Consider the decomposition in section 7.1
(50) $\xi=YA+QR.$
Let $\eta\in T_{Y}\mathrm{St}_{n,d}$ be another tangent vector at $X$. Then, the parallel transport $\Delta=\Delta(t)$ of $\eta$ along the geodesic $\gamma(t)$ with $\gamma(0)=Y,\dot{\gamma}(0)=\xi$ is given by
(51) $\begin{gathered}\Delta(t)=\begin{bmatrix}Y|Q\end{bmatrix}\operatorname{expm}(t\begin{bmatrix}2\alpha A&-R^{\mathsf{T}}\\ R&0\end{bmatrix})\operatorname{expv}(t\mathrm{P},\begin{bmatrix}Y|Q\end{bmatrix}^{\mathsf{T}}\eta)\operatorname{expm}(t(1-2\alpha)A)\\ +(\eta-YY^{\mathsf{T}}\eta-QQ^{\mathsf{T}}\eta)\operatorname{expm}(t(1-\alpha)A).\end{gathered}$
Here, $\mathrm{P}=\mathrm{P}_{AR}$ is the operator on $\mathcal{F}:=\mathrm{Skew}_{p}\times\mathbb{R}^{k\times p}\subset\mathbb{R}^{(p+k)\times p}$, mapping $w=\begin{bmatrix}w_{a}\\ w_{r}\end{bmatrix}\in\mathcal{F}$ to $\mathrm{P}(w)$ as follows
(52) $\mathrm{P}(w)=\mathrm{P}_{AR}(w):=\begin{bmatrix}((4\alpha-1)w_{a}A+R^{\mathsf{T}}w_{r})_{\mathrm{skew}}\\ \alpha(w_{r}A-Rw_{a})\end{bmatrix}.$
Proof.
Decompose $\eta=\eta_{\mathcal{S}}+\eta_{\mathcal{S}^{o}}$ with
$\begin{gathered}\eta_{\mathcal{S}}=YY^{\mathsf{T}}\eta+QQ^{\mathsf{T}}\eta,\\ \eta_{\mathcal{S}^{o}}=(\operatorname{I}_{n}-YY^{\mathsf{T}})\eta-QQ^{\mathsf{T}}\eta.\end{gathered}$
We can verify $\eta_{\mathcal{S}}$ is in $\mathcal{S}$ and $Y^{\mathsf{T}}\eta_{\mathcal{S}^{o}}=0$, $Q^{\mathsf{T}}\eta_{\mathcal{S}^{o}}=Q^{\mathsf{T}}\eta-Q^{\mathsf{T}}\eta=0$, hence $\xi^{\mathsf{T}}\eta_{\mathcal{S}^{o}}=0$. Thus, the theorem follows from eq. 49 applied to $\eta_{\mathcal{S}^{o}}$ and the isometry in section 7.1 applied to $\eta_{\mathcal{S}}$, and section 5. It remains to derive $\mathrm{P}$ from eq. 41. With $\mathtt{b}=\begin{bmatrix}\mathtt{b}_{A}&-\mathtt{b}_{R}^{\mathsf{T}}\\ \mathtt{b}_{R}&0\end{bmatrix}$, $\mathtt{a}=\begin{bmatrix}\mathtt{a}_{A}&-\mathtt{a}_{R}^{\mathsf{T}}\\ \mathtt{a}_{R}&0\end{bmatrix}$ then $[\mathtt{b},\mathtt{a}]_{\mathfrak{m}}\operatorname{I}_{d+k,d}=\begin{bmatrix}[\mathtt{b}_{A},\mathtt{a}_{A}]-\mathtt{b}_{R}^{\mathsf{T}}\mathtt{a}_{R}+\mathtt{a}_{R}^{\mathsf{T}}\mathtt{b}_{R}\\ \mathtt{b}_{R}\mathtt{a}_{A}-\mathtt{a}_{R}\mathtt{b}_{A}\end{bmatrix}$, $[\mathtt{a}_{\mathfrak{a}},\mathtt{b}]\operatorname{I}_{d+k,d}=\begin{bmatrix}[\mathtt{a}_{A},\mathtt{b}_{A}]\\ -\mathtt{b}_{R}\mathtt{a}_{A}\end{bmatrix}$, and $[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]\operatorname{I}_{d+k,d}=\begin{bmatrix}[\mathtt{b}_{A},\mathtt{a}_{A}]\\ -\mathtt{a}_{R}\mathtt{b}_{A}\end{bmatrix}$. Thus,
$\begin{gathered}\frac{1}{2}\left([\mathtt{b}_{A},\mathtt{a}_{A}]+\mathtt{a}_{R}^{\mathsf{T}}\mathtt{b}_{R}-\mathtt{b}_{R}^{\mathsf{T}}\mathtt{a}_{R}\right)-(1-2\alpha)[\mathtt{b}_{A},\mathtt{a}_{A}]=((4\alpha-1)\mathtt{b}_{A}\mathtt{a}_{A}+\mathtt{a}_{R}^{\mathsf{T}}\mathtt{b}_{R})_{\mathrm{skew}},\\ \frac{1}{2}(\mathtt{b}_{R}\mathtt{a}_{A}-\mathtt{a}_{R}\mathtt{b}_{A})+\frac{1-2\alpha}{2}(\mathtt{a}_{R}\mathtt{b}_{A}-\mathtt{b}_{R}\mathtt{a}_{A})=\alpha(\mathtt{b}_{R}\mathtt{a}_{A}-\mathtt{a}_{R}\mathtt{b}_{A}).\end{gathered}$
giving us the top and bottom rows of $\mathrm{P}_{\mathtt{a}}\mathtt{b}$, with $A=\mathtt{a}_{A},R=\mathtt{a}_{R},w_{r}=\mathtt{b}_{R},w_{a}=\mathtt{b}_{A}$. ∎
8. Flag manifolds
Flag manifolds find applications in data science in recent years, see [30, 27, 24, 17] for surveys. Additional background for flag manifolds could be found in [30, 21, 19]. The simplified parallel transport condition in section 6 is satisfied if $\beta=-1$, or $\alpha=\frac{1}{2}$, but not in general. We provide a summary relevant to the computation of geodesics and parallel transport.
With $n>d$, for a matrix $Y\in\mathrm{St}_{n,d}$, consider a partition of columns of $Y$ to $p$ blocks of sizes $(d_{1},\cdots,d_{p})$, $Y=\begin{bmatrix}Y_{1}|\cdots|Y_{p}\end{bmatrix}$, $Y_{i}\in\mathbb{R}^{n\times d_{i}}$. Consider the block-diagonal group $\mathrm{K}_{\mathfrak{a}}=\mathrm{S}(\operatorname{O}(d_{1})\times\cdots\times\operatorname{O}(d_{p}))\subset\mathrm{SO}(d)$, whose element $U=\operatorname{diag}(U_{1},\cdots U_{p})$ acts on $Y$ by right multiplication, $YU=\begin{bmatrix}Y_{1}U_{1}|\cdots|Y_{p}U_{p}\end{bmatrix}$ ($\mathrm{S}$ means $\det(U)=1$).
A flag manifold is the quotient of $\mathrm{St}_{n,d}$ by this action [30, section 4.2]. It is also typically expressed as a quotient of $\mathrm{SO}(n)$, allowing us to leverage our framework. Set $d_{p+1}=n-d$, the flag manifold $\mathrm{St}_{n,d}/\mathrm{K}_{\mathfrak{a}}$ will be denoted by $\operatorname{Flag}(\vec{d})$, where $\vec{d}=(d_{1},\cdots,d_{p},d_{p+1})$. We can also consider $\operatorname{Flag}(\vec{d})$ as a quotient of $\mathrm{SO}(n)$ by the diagonal blockdiagonal subgroup $\mathrm{K}=\mathrm{S}(\operatorname{O}(d_{1})\times\cdots\times\operatorname{O}(d_{p})\times\operatorname{O}(d_{p+1}))$. Going forward, we will use the term flag diagonal blocks to this block structure. Thus, in our setting, $\mathrm{G}=\mathrm{SO}(n),\mathfrak{g}=\mathfrak{so}(n),\mathfrak{a}=\mathfrak{so}(d)$ identified with the top left $d\times d$ block, $\mathfrak{k}=\mathfrak{so}(d_{1})\times\cdots\times\mathfrak{so}(d_{p})\times\mathfrak{so}(d_{p+1})$. Hence, $\mathfrak{k}$ is a block diagonal Lie algebra of antisymmetric matrices, while its complement $\mathfrak{m}$ is the space of antisymmetric matrices with zero flag diagonal blocks. $\mathfrak{k}_{\mathfrak{a}}$ is the Lie algebra of $\mathrm{K}_{\mathfrak{a}}$ above, while $\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}$ is the bottom $\mathfrak{so}(n-d)$. We show a diagram of the spaces in table 1 for $p=2$.
$n$ldelim{3* $\mathfrak{k}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$
$\mathfrak{m}_{\mathfrak{a}}$ $\mathfrak{k}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$
$\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$ $\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}=\mathfrak{a}_{{}_{\top}}$
$\scriptstyle d_{1}$ $\scriptstyle d_{2}$ $\scriptstyle d_{3}=n-d$
$\to$ $\mathfrak{k}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}}$ $\mathfrak{k}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$ $\scriptstyle d=d_{1}+d_{2}$
Table 1. Arrangement of subspaces and subalgebras for a flag manifold $\operatorname{Flag}(d_{1},d_{2},d_{3})$ as a quotient of $\mathrm{SO}(n)$ (left) or $\mathrm{St}_{n,d}$ (right) for $p=2$. The tangent space at $\operatorname{I}_{n}$, represented by an $n\times n$ matrix in $\mathfrak{so}(n)$ is divided to cells corresponding to subspaces and subalgebras. The top-left $2\times 2$ cells are in $\mathfrak{a}$, the bottom-right cell is $\mathfrak{a}_{{}_{\top}}$. The quotient is by the block-diagonal group $\mathrm{K}$ corresponding to $\mathfrak{k}=\mathfrak{k}_{\mathfrak{a}}\oplus\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}$. The metric is constructed using $\mathfrak{a}$. The image is the horizontal space at $\operatorname{I}_{n,d}\in\mathrm{St}_{n,d}$, with the last column block removed and the $\mathfrak{k}_{\mathfrak{a}}$ cells set to zero.
Under the metric with parameter $\vec{\beta}=(-\frac{1}{2},\alpha)$ in section 5, we note if $\mathtt{a}_{\mathfrak{a}}\in\mathfrak{a}$ with zero flag diagonal blocks (thus $\mathtt{a}_{\mathfrak{a}}\in\mathfrak{a}\cap\mathfrak{m}$), and $U(t)=\exp(t(2\alpha-1)\mathtt{a}_{\mathfrak{a}})$, then $U(t)^{\mathsf{T}}\mathfrak{k}U(t)$ may not be in $\mathfrak{k}$, and the simplified condition for parallel transport is not expected to apply in general. However, it applies for $\alpha=\frac{1}{2}$ where $U(t)=\operatorname{I}_{n}$. This is the bi-invariant metric on $\mathrm{SO}(n)$, the formula for $\mathrm{P}_{\mathtt{a}}$ in eq. 39 reduces to $\mathrm{P}_{\mathtt{a}}w=\frac{1}{2}[w,\mathtt{a}]_{\mathfrak{m}}$.
To compute in the more efficient Stiefel coordinates, we proceed as before. In Stiefel coordinates, the Levi-Civita connection is given by [21, eq. 6.4].
(53) $\nabla^{\mathcal{H}}_{\mathtt{Z}}\mathtt{W}=\mathrm{D}_{\mathtt{Z}}\mathtt{W}+Y(\mathtt{Z}^{\mathsf{T}}\mathtt{W})_{\mathrm{symf}}+(1-\alpha)(\operatorname{I}_{n}-YY^{\mathsf{T}})(\mathtt{Z}\mathtt{W}^{\mathsf{T}}+\mathtt{W}\mathtt{Z}^{\mathsf{T}})Y.$
Here, for a matrix $A\in\mathbb{R}^{d\times d}$, $A_{\mathrm{symf}}=\frac{1}{2}(A+A^{\mathsf{T}})+\frac{1}{2}(A-A^{\mathsf{T}})_{\mathfrak{k}_{\mathfrak{a}}}$, the projection to $\mathfrak{k}_{\mathfrak{a}}$ is the operation extracting the diagonal blocks corresponding to $\mathfrak{so}(d_{i})$, $i=1\cdots p$. Thus, we symmetrize $A$ but add back the antisymmetric parts of the flag diagonal blocks, hence $A_{\mathrm{symf}}$ means symmetrize $A$ but leaving the flag diagonal blocks unchanged. This accounts for the difference between the full tangent space of a Stiefel manifold and the horizontal space of a flag manifold, as seen from eq. 38. This is proved in [21]. Alternatively, we can project the formula eq. 38, similar to the Stiefel case in [20].
Thus, given an initial horizontal vector $\xi$ in $T_{Y}\mathrm{St}_{n,p}$, we can form the vector subspace $\mathcal{S}\subset\mathbb{R}^{n\times d}$ with columns spaned by columns of $Y$ and $\xi$. Decompose $\xi=YA+QR$ with $Q\in\mathbb{R}^{n\times k}$, then $\mathcal{M}_{\mathcal{S}}$ is isometric via the map $V\mapsto F_{YQ}V$ to $\mathrm{St}_{d+k,k}$, where $F_{YQ}=\begin{bmatrix}Y|Q\end{bmatrix}$ as before. The group $\mathrm{K}_{\mathfrak{a}}=\mathrm{S}(\operatorname{O}(d_{1})\times\cdots\times\operatorname{O}(d_{p}))$ operates on $\mathcal{M}_{\mathcal{S}}$ and on $\mathrm{St}_{d+k,k}$. By associativity $F_{YQ}(VU)=(F_{YQ}V)U$ if $U\in\mathrm{SO}(d)\supset\mathrm{K}_{\mathfrak{a}}$, thus horizontal spaces map to horizontal spaces. Therefore, flag geodesics and transports between these two spaces correspond, and we obtain geodesic formulas in reduced form. Similarly, we can map parallel transports with initial horizontal vectors in $\mathcal{M}_{\mathcal{S}}$, where we can use the simplified transport formula when $\alpha=\frac{1}{2}$.
More specifically, let $\eta$ be a horizontal vector to be transported, we decompose $\eta=\eta_{\mathcal{S}}+\eta_{\mathcal{S}^{o}}$, where $\eta_{\mathcal{S}^{o}}$ is horizontal and normal to both $Y$ and $\mathcal{S}$. Similar to the Stiefel case, the transport of $\eta_{\mathcal{S}^{o}}$ is given by eq. 49. This is proved exactly like in the Stiefel case, using eq. 53, with $\eta_{\mathcal{S}^{o}}=(\operatorname{I}_{n}-YY^{\mathsf{T}})\eta-QQ^{\mathsf{T}}\eta$, then $\gamma^{\mathsf{T}}\Delta_{\mathcal{S}^{o}}=\dot{\gamma}^{\mathsf{T}}\Delta_{\mathcal{S}^{o}}=0$, and the rest follows by solving a linear ODE explicitly. Thus,
Proposition 4.
For $\alpha=\frac{1}{2}$, at $Y\in\mathrm{St}_{n,p}$, with $\xi=YA+QR$ is a horizontal vector as in section 7, the parallel transport of a horizontal vector $\eta\in T_{Y}\mathrm{St}_{n,p}$ of $\operatorname{Flag}(\vec{d})$ lifted to $\mathrm{St}_{n,p}$ by
(54) $\begin{gathered}\Delta(t)=\begin{bmatrix}Y|Q\end{bmatrix}\operatorname{expm}(t\begin{bmatrix}A&-R^{\mathsf{T}}\\ R&0\end{bmatrix})\operatorname{expv}(t\mathrm{P},\begin{bmatrix}Y|Q\end{bmatrix}^{\mathsf{T}}\eta)\\ +(\eta-YY^{\mathsf{T}}\eta-QQ^{\mathsf{T}}\eta)\operatorname{expm}(\frac{t}{2}A).\end{gathered}$
With $\mathtt{a}=\begin{bmatrix}A&-R^{\mathsf{T}}\\ R&0\end{bmatrix}$ and $w=\begin{bmatrix}w_{a}&-w_{r}^{\mathsf{T}}\\ w_{r}&0\end{bmatrix}$ then $\mathrm{P}$ is given by
(55) $\mathrm{P}(w)=\mathrm{P}_{\mathtt{a}}(w):=\begin{bmatrix}((w_{a}A+R^{\mathsf{T}}w_{r})_{\mathrm{skew}})_{\mathfrak{m}}\\ \frac{1}{2}(w_{r}A-Rw_{a})\end{bmatrix}.$
For the proof, we simplify eq. 52. The subscript $\mathfrak{m}$ on the top $d\times d$ block means setting the flag diagonal blocks to zero.
9. Numerical studies
We implemented the formulas in Python. The tables and graphs are produced from Google colab workbooks in the folder examples in [22] on default free session setting with two Xenon CPUs and $13$GB of RAM.
As mentioned, the algorithm [2] for exponential action requires a norm estimate for the operator $\mathrm{P}$. For Stiefel and flag manifolds, after a rescaling, this operator becomes antisymmetric in the Frobenius norm. Numerically, in this case, the estimate using just the 1-norm $\|\mathrm{P}\|_{1}$ is adequate, instead of a higher $p$-norm estimate.
Our next observation is for our problem, $\|\mathrm{P}\|_{1}$ can be estimated more efficiently than by calling a library function. Recall that on a $n$-dimension vector space with a basis $\{e_{i}\}_{i=1}^{n}$, the 1-norm of a vector $\sum c_{i}e_{i}$ is $\sum_{i}|c_{i}|$. Vectorizing an operator $\mathrm{P}$ using this basis, we have $\|\mathrm{P}\|_{1}=\max_{j}|\mathrm{P}e_{j}|_{1}$, which has a cost of $O(n^{2})$, thus, it is $O(m^{2}d^{2})$ on a matrix space $\mathbb{R}^{m\times d}$. The sampling in [10] reduces this cost, but we can reduce it further by exploiting the structure of $\mathrm{P}$. For the Stiefel and flag manifold cases, we provide an estimate based on norms of $A$ and $R$. The resulting customized implementation of exponential action could be more than $10$ times faster than the library version, significantly reducing the total time cost of the parallel transport calculation. To keep focus, we do not attempt these estimates of $\|\mathrm{P}\|_{1}$ for $\operatorname{GL}^{+}(n)$ and $\mathrm{SO}(n)$, while they seem likely to be available. Likewise, we will not discuss the scalings to make $\mathrm{P}$ antisymmetric for these manifolds as in lemma 4 below.
Further, the scipy function $expm{\_}multiply$ is not yet implemented in JAX [4], a Python library that provides easy compilation to run in GPU and allows automatic differentiation (AD). We use our estimates to implement a simple version of $expv$ in JAX, and use the AD feature to compute the derivative of the parallel transport, verifying that our implementation is highly accurate. We offer $expm{\_}multiply$-based parallel transport for all four manifolds while providing customized implementations for Stiefel and flag manifolds, with AD(JAX) and without-AD versions. The discussion here will likely apply to other languages using similar algorithms.
First, for a real number $c$, define $s_{c}$ to be the operator on $\mathbb{R}^{m\times d}$ with $m\geq d$, scaling the top $d\times d$ block by $c$ and preserving the bottom $(m-d)\times d$ block. We have $\operatorname{expm}\mathrm{P}=s_{c}^{-1}\operatorname{expm}(s_{c}\mathrm{P}s_{c}^{-1})s_{c}$ for an operator $\mathrm{P}$ on $\mathbb{R}^{m\times d}$. The following computes an upper bound for $\|s_{\alpha^{\frac{1}{2}}}\mathrm{P}_{AR}s_{\alpha^{-\frac{1}{2}}}\|_{1}$ with a cost of $O(d^{2})$:
Lemma 4.
For $k\geq 0$ and $A=(a_{ij})_{ij=(1,1)}^{(d,d)}\in\mathfrak{so}(d),R=(r_{ij})_{ij=(1,1)}^{(k,d)}\in\mathbb{R}^{k\times d}$, with respect to the standard basis of $\mathbb{R}^{(d+k)\times d}$, the operator $\mathrm{P}_{AR}^{bal}:=s_{\alpha^{\frac{1}{2}}}\circ\mathrm{P}_{AR}\circ s_{\alpha^{-\frac{1}{2}}}$ with $\mathrm{P}_{AR}$ in eq. 52 operating on $\mathbb{R}^{(d+k)\times d}$ has an upper bound on the $1$-norm
(56) $\begin{gathered}\|\mathrm{P}_{AR}^{bal}\|_{1}=\max_{|w|_{1}=1}\lvert\mathrm{P}_{AR}^{bal}w\rvert_{1}\leq\max(n_{A},n_{R})\\ \text{where }n_{A}=\max_{j}(\alpha^{\frac{1}{2}}\sum_{i=1}^{d}(\lvert r_{ij}\rvert+\lvert 4\alpha-1\rvert\|A\|_{1})),\\ n_{R}=\max_{j}(\alpha\sum_{i}(\lvert a_{ij}\rvert+\alpha^{\frac{1}{2}}\|R\|_{\infty})).\end{gathered}$
Also, define the operator $\mathsf{skew}_{top}$ on $\mathbb{R}^{(d+k)\times d}$ to be the map antisymmetrizing the top $d\times d$ block, then $\mathrm{P}_{AR}^{bal}\circ\mathsf{skew}_{top}:\begin{bmatrix}w_{a}\\ w_{r}\end{bmatrix}\mapsto\mathrm{P}_{AR}^{bal}\begin{bmatrix}(w_{a})_{\mathrm{skew}}\\ w_{r}\end{bmatrix}$ is antisymmetric in the Frobenius inner product. With $\mathrm{P}^{\operatorname{Flag}}_{AR}$ given in eq. 55 at $\alpha=\frac{1}{2}$, we have the same bound for $s_{\alpha^{\frac{1}{2}}}\circ\mathrm{P}^{\operatorname{Flag}}_{AR}\circ s_{\alpha^{-\frac{1}{2}}}$.
Proof.
For $w=\begin{bmatrix}w_{a}\\ w_{r}\end{bmatrix}$, we have
$\mathrm{P}_{AR}^{bal}w=\begin{bmatrix}\alpha^{\frac{1}{2}}((4\alpha-1)\alpha^{-\frac{1}{2}}w_{a}A+R^{\mathsf{T}}w_{r})_{\mathrm{skew}}\\ \alpha(w_{r}A-\alpha^{-\frac{1}{2}}Rw_{a})\end{bmatrix}=\begin{bmatrix}((4\alpha-1)w_{a}A+\alpha^{\frac{1}{2}}R^{\mathsf{T}}w_{r})_{\mathrm{skew}}\\ \alpha w_{r}A-\alpha^{\frac{1}{2}}Rw_{a}\end{bmatrix}.$
Consider the basis vectors $E_{ij}$ with $1\leq i\leq d+k,1\leq j\leq d$, where $E_{ij}$ is an elementary matrix with the $(i,j)$-entry is one and other entries vanish. Write $A=\sum_{qs=(1,1)}^{(d,d)}a_{qs}E_{qs},R=\sum_{qs=(1,1)}^{(k,d)}r_{qs}E_{(d+q)s}$. For $i\leq d$,
$\begin{gathered}|\mathrm{P}^{bal}_{AR}E_{ij}|_{\vec{1}}=\bigg{\lvert}\begin{bmatrix}\frac{4\alpha-1}{2}(\sum_{r=1}^{d}a_{jr}E_{ij}E_{jr}-\sum_{r=1}^{d}a_{jr}E_{rj}E_{ji})\\ -\alpha^{\frac{1}{2}}\sum_{q=1}^{k}r_{qi}E_{(d+q)i}E_{ij}\end{bmatrix}\bigg{\rvert}_{\vec{1}}\\ \leq\frac{|4\alpha-1|}{2}(\sum_{r=1}^{d}|a_{jr}|+\sum_{s=1}^{d}|a_{js}|)+\alpha^{\frac{1}{2}}\sum_{q=1}^{k}|r_{qi}|\leq n_{A}\end{gathered}$
where |.$|.|_{\vec{1}}$ denotes the $1$-norm after vectorization, because $E_{ij}E_{jr}=E_{ir}$ is distinct from $E_{sj}E_{ji}=E_{si}$ unless $i=r=s$, in this case, we have a cancelation. For $i>d$
$\begin{gathered}|\mathrm{P}^{bal}_{AR}E_{ij}|_{\vec{1}}=\bigg{\lvert}\begin{bmatrix}\alpha^{\frac{1}{2}}(\sum_{s=1}^{d}r_{i-d,s}E_{si}E_{ij})_{\mathrm{skew}}\\ \alpha E_{ij}\sum_{s}a_{js}E_{js}\end{bmatrix}\bigg{\rvert}_{\vec{1}}\leq\frac{2\alpha^{\frac{1}{2}}}{2}\sum_{s}|r_{i-d,s}|+\alpha\sum_{s}|a_{js}|\leq n_{R}.\end{gathered}$
This gives us the bound. To see $\mathrm{P}_{AR}^{bal}\circ\mathsf{skew}_{top}$ is antisymmetric, note the diagonal parts $w_{a}\mapsto(4\alpha-1)((w_{a})_{\mathrm{skew}}A)_{\mathrm{skew}}$ and $w_{r}\mapsto\alpha w_{r}A$ are both antisymmetric, since $A$ is antisymmetric. We can verify directly the off-diagonal operators $F_{1}:w_{r}\mapsto\alpha^{\frac{1}{2}}(R^{\mathsf{T}}w_{r})_{\mathrm{skew}}$ and $F_{2}:w_{a}\mapsto-\alpha^{\frac{1}{2}}R(w_{a})_{\mathrm{skew}}$ satisfy $F_{1}=-F_{2}^{\mathsf{T}}$. For flag manifolds, clearing flag diagonal blocks does not increase norms, thus the same estimate holds. ∎
With this estimate, following scipy’s implementation, we have a finer extension of [2, table 3.1] containing a look up table for constants $\theta_{m}$ from $m$. The number of Taylor terms $m_{*}$ will be the value of $m$ minimizing $\mathrm{ceil}(\frac{m\|\mathrm{P}\|_{1}}{\theta_{m}})$ and the scaling parameter $s$ is set to be $\mathrm{ceil}(\frac{\|\mathrm{P}\|_{1}}{\theta_{m_{*}}})$, both are used in the approximation $\operatorname{expv}(\mathrm{P},\mathtt{b})\approx(T_{m}(s^{-1}\mathrm{P}))^{s}\mathtt{b}$.
In table 2, we show the computational time along a time grid, for different $n$ and $\alpha$ for Stiefel manifolds, and for $\alpha=\frac{1}{2}$ (the canonical metric). The flag manifold implementation simplifies the transport formula when $\alpha=\frac{1}{2}$, resulting in slightly faster execution. Table table 3 shows execution time by $d$, also along the same time grid. Using automatic differentiation, in [22] we can verify that these correspond to highly accurate computation, depending on the manifold, with the transport equation satisfied with accuracy around $10^{-10}$ to $10^{-12}$.
Stiefel man. comp. time (s) Flag man. comp. time.(s)
Time grid t(s) for geodesics Time grid t(s) for geodesics
$\alpha$ n 0.5 1.0 2.0 5.0 20.0 0.5 1.0 2.0 5.0 20.0
$\frac{1}{2}$ 100 0.04 0.08 0.14 0.25 0.59 0.04 0.07 0.12 0.17 0.50
200 0.07 0.13 0.16 0.25 0.75 0.06 0.10 0.16 0.23 0.56
1000 0.13 0.19 0.25 0.48 1.23 0.11 0.17 0.21 0.34 1.33
2000 0.23 0.28 0.35 0.61 1.80 0.23 0.27 0.32 0.54 1.60
1 100 0.08 0.12 0.17 0.32 1.13
200 0.09 0.14 0.20 0.37 1.44
1000 0.17 0.21 0.33 0.61 2.10
2000 0.26 0.32 0.46 0.76 2.73
Table 2. Computational time for parallel transport for Stiefel and flag manifolds by $n$ for $d=50$. The flags are $(20,15,15,n-d)$. The canonical metric corresponds to $\alpha=\frac{1}{2}$, the Euclidean embedded metric corresponds to $\alpha=1$. The table shows approximate linear time in $n$ and $t$. Execution is on a free colab session.
Stiefel man. comp. time (s) Flag man. comp. time.(s)
Geo. time grid t(s) Geo. time grid t(s)
$\alpha$ d 0.5 1.0 2.0 5.0 20.0 0.5 1.0 2.0 5.0 20.0
$\frac{1}{2}$ 5 0.03 0.04 0.06 0.13 0.29
10 0.03 0.05 0.08 0.15 0.40 0.03 0.05 0.10 0.15 0.37
20 0.06 0.08 0.14 0.19 0.45 0.06 0.08 0.12 0.17 0.41
100 0.35 0.59 0.97 2.15 8.39 0.30 0.48 0.70 2.00 7.19
1 5 0.04 0.07 0.11 0.15 0.40
10 0.04 0.06 0.09 0.17 0.61
20 0.07 0.10 0.16 0.26 0.66
100 0.54 0.76 1.54 3.51 14.07
Table 3. Computational time for parallel transport for Stiefel and flag manifolds by $d$ for $n=1000$. We scale the partial flag of $(d_{1},d_{2},d_{3})=(4,3,3)$ by $\frac{d}{4+3+3}$ to $d$ for flag manifolds, setting $d_{4}=n-d$. Computational time is nonlinear in $d$.
9.1. Parallel transport as isometries
We also demonstrate numerically the isometric property of parallel transport as implemented by our method, by showing the inner product matrix of a set of vectors is preserved during transportation. We consider a geodesic starting at $\operatorname{I}_{2000,100}$, with twenty tangent vectors having integer lengths sampled from integers between $1$ and $60$, while the relative angles are randomly generated. The geodesic initial velocity vector $v$ has length $1$. We compute the inner product matrix of the random tangent vectors at $t=0$ and of the transported vectors along a time grid. We measure the maximum absolute difference between the inner product matrices at each grid point and the initial point and plot the $\log_{10}$ of the differences in fig. 1. The absolute errors are around $10^{-12}$ up to $t=15$.
(a) Stiefel manifolds
(b) Flag manifolds
[…]
10. Discussion
In this paper, we provide an efficient algorithm to compute parallel transport on Stiefel and flag manifolds, as well as transport formulas for the generalized linear group and the special orthogonal group. We also provide a framework for parallel transports on groups with transposable Lie algebra. The symplectic group [6] also has transposable Lie algebra, thus the framework here should apply. We believe the framework developed here will be useful for other manifolds in applied mathematics.
Acknowledgments
Du Nguyen would like to thank his family for their support in this project. Stefan Sommer is supported by a research grant (VIL40582) from VILLUM FONDEN.
Appendix A The Grassmann case
The Grassmann manifold is a flag manifold with $p=1$. In this case, with $\vec{d}=(d,n-d)$, $\alpha=\frac{1}{2}$, we have $\mathrm{P}$ is identically zero, $A=0$, $Y^{\mathsf{T}}\eta=0$, thus, $\begin{bmatrix}Y|Q\end{bmatrix}^{\mathsf{T}}QQ^{\mathsf{T}}\eta=\begin{bmatrix}0\\ Q^{\mathsf{T}}\end{bmatrix}\eta$. We have
$\Delta(t)=\begin{bmatrix}Y|Q\end{bmatrix}\operatorname{expm}(t\begin{bmatrix}0&-R^{\mathsf{T}}\\ R&0\end{bmatrix})\begin{bmatrix}0\\ Q^{\mathsf{T}}\end{bmatrix}\eta+\eta-QQ^{\mathsf{T}}\eta.$
If $\eta=Q\Sigma V^{\mathsf{T}}=QR$ is a compact SVD decomposition, where $R:=\Sigma V^{\mathsf{T}}$, with $\Sigma$ is diagonal in $\mathbb{R}^{k\times k}$, $V\in\mathbb{R}^{k\times d}$ with $V^{\mathsf{T}}V=\operatorname{I}_{d}$. Set $C=\begin{bmatrix}V&0\\ 0&\operatorname{I}_{d}\end{bmatrix}\in\mathbb{R}^{(d+k)\times(d+k)}$, then
$\begin{gathered}\begin{bmatrix}0&-R^{\mathsf{T}}\\ R&0\end{bmatrix}=C\begin{bmatrix}0&-\Sigma\\ \Sigma&0\end{bmatrix}C^{\mathsf{T}},\\ \operatorname{expm}(t\begin{bmatrix}0&-R^{\mathsf{T}}\\ R&0\end{bmatrix})=\begin{bmatrix}\operatorname{I}_{d}-VV^{\mathsf{T}}&0\\ 0&0\end{bmatrix}+C\begin{bmatrix}\cos(t\Sigma)&-\sin(t\Sigma)\\ \sin(t\Sigma)&\cos(t\Sigma)\end{bmatrix}C^{\mathsf{T}},\end{gathered}$
where we use $\operatorname{C}^{\mathsf{T}}\operatorname{C}=\operatorname{I}_{d+k}$ to simplify the middle terms $\frac{t^{j}}{j!}(C\begin{bmatrix}0&-\Sigma\\ \Sigma&0\end{bmatrix}C^{\mathsf{T}})^{j}$, $j>0$ in the power series expansion of $\operatorname{expm}$, while the constant term has the adjustment with the term $\operatorname{I}_{d}-VV^{\mathsf{T}}$ above. Thus, $\operatorname{expm}(t\begin{bmatrix}0&-R^{\mathsf{T}}\\ R&0\end{bmatrix})\begin{bmatrix}0\\ Q^{\mathsf{T}}\end{bmatrix}\eta$ simplifies to $C\begin{bmatrix}-\sin(t\Sigma)\\ \cos(t\Sigma)\end{bmatrix}Q^{\mathsf{T}}\eta$, and
(57) $\Delta(t)=\begin{bmatrix}YV|Q\end{bmatrix}\begin{bmatrix}-\sin(t\Sigma)\\ \cos(t\Sigma)\end{bmatrix}Q^{\mathsf{T}}\eta+\eta-QQ^{\mathsf{T}}\eta,$
which is eq (2.68) in [5].